### v1

# EarthquakeNPP: Benchmark Datasets for Earthquake Forecasting with Neural Point Processes

Abstract
Classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes for decades. Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over classical models. However, the currently-used benchmark dataset for NPPs does not represent an up-to-date challenge in the seismological community since it lacks a key earthquake sequence from the region and improperly splits training and testing data. Furthermore, initial earthquake forecast benchmarking lacks a comparison to state-of-the-art earthquake forecasting models typically used by the seismological community. To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by a credible implementation of the ETAS model. The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation. In a benchmarking experiment, we compare three spatio-temporal NPPs against ETAS and find that none outperform ETAS in either spatial or temporal log-likelihood. These results indicate that current NPP implementations are not yet suitable for practical earthquake forecasting. However, EarthquakeNPP will serve as a platform for collaboration between the seismology and machine learning communities with the goal of improving earthquake predictability.
1 Introduction
Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes. While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model (Ogata, 1988; 1998), despite significant growth in available data (Takanami et al., 2003; Shelly, 2017; Ross et al., 2019; White et al., 2019; Mousavi et al., 2020; Tan et al., 2021; Mousavi & Beroza, 2023).
In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019; Jia & Benson, 2019; Chen et al., 2021; Zhou et al., 2022; Zhou & Yu, 2024). While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community. The benchmark lacks a key earthquake sequence from the region, fails to recreate an operational setting with proper train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
Here, we introduce EarthquakeNPP: a curated collection of datasets designed for benchmarking NPP models in earthquake forecasting, accompanied by a state-of-the-art benchmark model. These datasets are derived from publicly available raw data, which we process and configure within our platform to facilitate meaningful forecasting experiments relevant to stakeholders in the seismology community. Covering various regions of California, these datasets represent typical forecasting zones and encompass data commonly utilized by forecast issuers. Moreover, employing modern techniques, some datasets include smaller magnitude earthquakes, exploring the potential of numerous small events to enhance forecasting performance through flexible NPPs. To unify efforts, we present an operational-level implementation of the ETAS model alongside the datasets, serving as a benchmark for NPPs.
Although initial benchmarking finds that none of the 3 tested NPP implementations outperform ETAS, EarthquakeNPP aims to serve as a platform for future NPP development. The platform facilitates the generative evaluation procedure used for rigorous benchmarking in the seismology community, directing the impact of future NPPs to stakeholders in seismology. Access to the dataset collection, along with comprehensive documentation and notebooks, can be found at https://anonymous.4open.science/r/EarthquakeNPP-2D51.
1.1 Related Work
Existing Benchmark Dataset. Chen et al. (2021) introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S. Geological Survey, focusing on Japan from 1990 to 2020. They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset. They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake (Mori et al., 2011), the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0. This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact. Additionally, these events are of significant scientific interest because they provide valuable insights into the earthquake rupture process.
The dataset segments are divided for training, testing, and validation. Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern. This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering (Freed, 2005). Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards it time.
Although earthquakes with magnitudes above 2.5 are considered by Chen et al. (2021), following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset. For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets. Section A.2 describes the biases incurred from such data missingness.
Chen et al. (2021) benchmark their model against another spatio-temporal model, Neural Jump SDEs (Jia & Benson, 2019), and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark. Subsequent papers adopting this benchmark (Zhou et al., 2022; Yuan et al., 2023; Zhou & Yu, 2024) similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
Temporal-NPP Benchmarking on Earthquake Data. Two existing works benchmark NPPs for earthquake forecasting within the seismology community. The first by Dascher-Cousineau et al. (2023) extends a temporal-only NPP from Shchur et al. (2019) to include earthquake magnitudes. The second by Stockman et al. (2023) extends another temporal-only model by Omi et al. (2019a) to target larger magnitude events. Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline. Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
Benchmarking within the Seismology Community. Model comparison has been crucial in the development of earthquake forecasting models since their inception (Kagan & Knopoff, 1987; Ogata, 1988). The Collaboratory for the Study of Earthquake Predictability (CSEP) (Michael & Werner, 2018; Schorlemmer et al., 2018; Savran et al., 2022; Iturrieta et al., 2024) (https://cseptesting.org/ ) aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally. CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day). These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times. The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
1.2 Scope of this work
Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development (Ogata, 1988; Harte, 2015). Our platform uses this metric in the NPP benchmarking experiment and provides detailed guidance on CSEP’s simulation-based procedure, enabling future NPP implementations and evaluations within CSEP experiments.
The goal of this work is to allow Machine Learning researchers to have seismological impact by defining a baseline target for which NPP models can be compared to state-of-the-art domain-based models. NPPs that can generate log-likelihoods are in scope, whilst those that do not (e.g. Yuan et al., 2023; Li et al., 2023) are out of scope because a valid score does not currently exist. The popular next-event point prediction metrics (e.g. Root Mean Square Error (RMSE) and related scores) are considered to be flawed and misleading for seismological prediction (Hodson, 2022), because the predictive distribution is strongly skewed and therefore far from Gaussian. To have seismological relevance, authors of NPP models are challenged to implement long-term predictions using CSEP’s evaluation procedure, benchmarking against the reported performance for the ETAS model.
2 Background
2.1 Spatio-Temporal Point Processes
A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\in\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$. This process is typically defined by a non-negative conditional intensity function
$\lambda(t,\mathbf{x}|\mathcal{H}_{t}):=\lim_{\Delta t,\Delta\mathbf{x}\rightarrow 0}\frac{\mathbb{E}\left[N([t,t+\Delta t)\times B(\mathbf{x},\Delta\mathbf{x})|\mathcal{H}_{t}\right]}{|B(\mathbf{x},\Delta\mathbf{x})|},$ (1)
where $\mathcal{H}_{t}=\{(t_{i},\mathbf{x}_{i})|t_{i}<t\}$ denotes the history of events preceding time $t$ and $|B(\mathbf{x},\Delta\mathbf{x})|$ is the Lebesgue measure of the ball $B(\mathbf{x},\Delta\mathbf{x})$ with radius $\Delta\mathbf{x}$. Given we observe a history of events up to $t_{i}$, the probability density function (pdf) of observing an event at time $t$ and location $\mathbf{x}$ is given by
[…]
Most models specify the conditional intensity function, though some (e.g. Shchur et al., 2019; Chen et al., 2021; Yuan et al., 2023), directly model this pdf. Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $[T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
[…]
where the decomposition of the spatio-temporal conditional intensity function, $\lambda(t_{i},\mathbf{x}_{i}|\mathcal{H}_{t_{i}})=\lambda(t_{i}|\mathcal{H}_{t_{i}})\cdot f(\mathbf{x}_{i}|t_{i},\mathcal{H}_{t_{i}})$, allows the log-likelihood to be written as contributions from the temporal and spatial components. In practice, this exact function is often not maximized directly during training: for models specified through the conditional intensity function, an analytical solution to the integral term is generally not possible and is approximated numerically.
For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric. This is consistent with a wealth of literature in the seismology community (see Zechar et al., 2010, and references therein) as well as the wider general point process literature (Daley & Vere-Jones, 2004), which now more recently includes neural point processes (Shchur et al., 2021). The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident. Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $[T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function. Comparing the mean log-likelihood per event provides the information gain from one model to another (Daley & Vere-Jones, 2004).
2.2 ETAS
The Epidemic Type Aftershock Sequence (ETAS) model (Ogata, 1998) is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space. It has been adopted for operational earthquake forecasting by government agencies in California (Milner et al., 2020), New-Zealand (Christophersen et al., 2017), Italy (Spassiani et al., 2023), Japan (Omi et al., 2019b) and Switzerland (Mizrahi et al., 2024), and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments (e.g. Woessner et al., 2011; Rhoades et al., 2018; Taroni et al., 2018; Cattania et al., 2018; Mancini et al., 2019; 2020; 2022). The general formulation of the model is
$\lambda(t,\mathbf{x}|\mathcal{H}_{t};\mathbf{\theta})=\mu+\sum_{i:t_{i}<t}g(t-t_{i},||\mathbf{x}-\mathbf{x}_{i}||^{2}_{2},m_{i}),$ (4)
where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event. The equivalent formulation as a Hawkes branching process accompanies a causal branching structure $\mathbf{B}$. This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g. via dynamic wave triggering (Brodsky & van der Elst, 2014) and static stress triggering (Gomberg, 2018; Mancini et al., 2020).
[…]
$\mathbf{\theta}^{(k+1)}=\operatorname*{arg\,max}_{\mathbf{\theta}}\mathbb{E}_{\mathbf{B}\sim p(\cdot|\mathcal{H}_{T_{1}},\mathbf{\theta}^{(k)})}\left[\log p(\mathcal{H}_{T_{1}},\mathbf{B}|\mathbf{\theta})\right].$ (5)
This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation and simultaneously estimates the causal structure.
The formulation of the ETAS model we present with the EarthquakeNPP datasets is implemented in the etas python package by Mizrahi et al. (2022). It defines the triggering kernel as
$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{cut})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{cut})}\right)^{1+\rho}},$ (6)
where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$.
3 EarthquakeNPP Datasets
The EarthquakeNPP datasets encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021. California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms (Gerstenberger et al., 2004; Field, 2007; Field et al., 2021). It encompasses the San Andreas fault plate boundary system (Zoback et al., 1987) and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
Notebooks to access and preprocess these public datasets along with the associated benchmarking experiment are publicly accessible at https://anonymous.4open.science/r/EarthquakeNPP-2D51, accompanied by more detailed documentation for each dataset. A summary of how earthquake datasets are generated, along with the associated challenges of using earthquake catalog data can be found in Appendix A. The following subsections provide a short overview of each EarthquakeNPP dataset.
3.1 ANSS Comprehensive Earthquake Catalog (ComCat)
The U.S. Geological Survey (USGS) National Earthquake Information Center (NEIC) monitors global earthquakes (Mw 4.5 or larger) and provides complete seismic monitoring of the United States for all significant earthquakes (> Mw 3.0 or felt). Its contributing seismic networks have produced the Advanced National Seismic System (ANSS) Comprehensive Catalog of Earthquake Events and Products. We focus on the California region defined by Schorlemmer & Gerstenberger (2007), with a test period consistent with CSEP experiments (Zechar et al., 2013).
3.2 Southern California Earthquake Data Center (SCEDC) Catalog
The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California (Hutton et al., 2010) since the Caltech Seismological Laboratory began routine operations in 1932. Significant network improvements since the 1970s and 1980s reduced the catalog completeness from Mw 3.25 to Mw 1.8. We use three magnitude thresholds (Mw 2.0, 2.5, 3.0) to explore the effect of truncation on forecasting model performance. Training includes the Mw 7.3 Landers and 1999 Mw 7.1 Hector Mine earthquakes, while testing involves the 2019 Mw 7.1 Ridgecrest sequence.The USGS utilizes both ComCat and SCEDC datasets in the aftershock forecasts they release to the public. The inclusion of these datasets determines whether NPPs can exploit the datasets currently being used for operational forecasting.
3.3 Detailed Earthquake Catalog for the San Jacinto Fault-Zone Region
White et al. (2019) created an enhanced catalog focusing on the San Jacinto fault region, using an dense seismic network in Southern California. This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion (White et al., 2019). The training window includes the 2010 Mw 5.4 Borrego Springs and 2013 ML 4.7 Anza Borrego earthquakes. This catalog is named White after the authors.
Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP. Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans. In red is a fault map from the GEM Global Active Faults Database (Styron & Pagani, 2020).
3.4 Quake Template Matching (QTM) Catalog
Using data collected by the SCSN, Ross et al. (2019) generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events. The catalog contains 1.81 million earthquakes complete down to Mw 0.3. Following Dascher-Cousineau et al. (2023), we use a more conservative completeness estimate of Mw 1.0 and split the catalog into two focus regions: the San Jacinto fault region and the Salton Sea. The inclusion of White, QTM_SanJac and QTM_SaltonSea determines whether very low magnitude earthquakes exhibit different behaviour to those of the larger earthquakes, an assumption which ETAS makes.
3.5 Additional Datasets
Beyond the official EarthquakeNPP datasets, we include 3 further datasets that either provide additional scientific insight or continuity from previous benchmarking works.
[…]
$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (7)
where $M$ is the mainshock magnitude. Events below this threshold are removed using mainshocks of Mw 5.2 and above. The inclusion of these datasets allows us to test whether NPPs are inhibited by data missingness to the same extent that ETAS is.
Deprecated Catalog of Japan. To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al. (2021), however with a chronological train-test split and without removing any supposed outlier events. To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
4 Benchmarking Experiment
We now use EarthquakeNPP to benchmark three spatio-temporal NPPs with prior positive claims on earthquake forecasting.
Neural Spatio-Temporal Point Process (NSTPP) (Chen et al., 2021): a pdf based NPP that parameterizes the spatial pdf with continuous-time normalizing flows (CNFs). We use their Attentive CNF model for its computational efficiency and overall performance versus their other model Jump CNF (Chen et al., 2021).
Deep Spatio-Temporal Point Process (Deep-STPP) (Zhou et al., 2022): a conditional intensity function based NPP that constructs a non parametric space-time intensity function governed by a deep latent process. The intensity function enjoys a closed form integration, avoiding the need for numerical approximation.
Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) (Zhou & Yu, 2024): a conditional intensity function based NPP which jointly models the 3D space-time integral of the intensity along with its derivative (the intensity function) using a dual network approach.
The benchmark is against the ETAS model defined in section 2.2, as well as a homogeneous Poisson process. The Poisson model is fit to events in the auxiliary, training and validation windows to provide a baseline score against which to compare all four other models.
Validation is typically not part of the estimation procedure for ETAS, so it is fit using the combined training and validation windows. NPPs follow the standard training/validation/testing procedure of machine learning. When possible, a model’s likelihood for training, validation, and testing can depend on events occurring before the splits through memory in its history. The exception is NSTPP, lacking a direct dependency on prior events. Nonetheless, its likelihood is evaluated on the same events as the other models. The definition of the ETAS model (equation 4) specifies how the magnitudes of earthquakes in the history contribute towards the intensity function. This earthquake magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modeling choices beyond the scope of this work.
Figures 2 and 3 report the temporal and spatial log-likelihood scores of all models on the EarthquakeNPP datasets. The ETAS model achieves the highest temporal and spatial log-likelihood across all datasets, with some NPP models achieving comparable temporal performance on ComCat, QTM_SaltonSea, QTM_SanJac, and White catalogs. Amongst the NPP models, Deep-STPP generally performs best in terms of temporal log-likelihood, whereas AutoSTPP performs best in terms of spatial log-likelihood. The improved relative temporal performance of all NPPs compared to ETAS as the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, as well as the comparable performance to ETAS using the low magnitude catalogs QTM_SaltonSea, QTM_SanJac, and White, indicates that low magnitude earthquakes provide valuable predictive information for NPPs. Appendix B provides further results using the additional datasets.
Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Figure 3: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
5 CSEP Consistency Tests
EarthquakeNPP supports the earthquake forecast evaluation protocol developed by the Collaboratory for the Study of Earthquake Predictability (CSEP). In this procedure a model generates 24-hour forecasts through $10,000$ repeat simulations of earthquake sequences at the beginning of every day in the testing period. This procedure exactly mimics how earthquake forecasts are generated in an operational setting (van der Elst et al., 2022). Models can then be evaluated by comparing the observed sequence with the distribution over model simulations. Three test statistics target the temporal, spatial and magnitude components of the forecasts, where a test is failed if the observed statistic falls within a pre-defined rejection region (Figure 4). We demonstrate this procedure for the ETAS model and report performance scores as a benchmark for future immplementations of NPPs.
5.1 Number Test
The number test evaluates the temporal component of the forecast by checking the consistency of the forecasted number of events, $N$ with those observed in the forecast horizon, $N_{\text{obs}}$. Upper and lower quantiles are estimated using the empirical cumulative distribution from the repeat simulations, $F_{N}$,
$\delta_{1}$ $=\mathbb{P}(N\geq N_{\text{obs}})=1-F_{N}(N_{\text{obs}}-1)$ (8)
$\delta_{2}$ $=\mathbb{P}(N\leq N_{\text{obs}})=F_{N}(N_{\text{obs}}).$ (9)
5.2 Spatial Test
To evaluate the spatial component of the forecast, a test statistic aggregates the forecasted rates of earthquakes over a regular grid,
$S=\left[\sum_{i=1}^{N}\log\hat{\lambda}(k_{i})\right]N^{-1},$ (10)
where $\hat{\lambda}(k_{i})$ is the approximate rate in the cell $k$ where the $i^{th}$ event is located. Upper and lower quantiles are estimated by comparing the observed statistic
$S_{\text{obs}}=\left[\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}(k_{i})\right]N_{\text{obs}}^{-1},$ (11)
with the empirical cumulative distribution of $S$ using the repeat simulations, $F_{S}$
$\gamma_{s}=\mathbb{P}(S\leq S_{\text{obs}})=F_{S}(S_{\text{obs}}).$ (12)
The grid is constructed from $\{0.1\degree,0.05\degree,0.01\degree\}$ squares for ComCat, SCEDC and {QTM_Salton_Sea, QTM_SanJac, White} respectively.
5.3 Magnitude Test
To evaluate the earthquake magnitude component of the forecast, a test statistic compares the histogram of a forecast’s magnitudes $\Lambda^{(m)}$, against the mean histogram over all forecasts $\bar{\Lambda}^{(m)}$,
$D=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda^{(m)}(k)+1\right]\right)^{2},$ (13)
where $\Lambda^{(m)}(k)$ and $\bar{\Lambda}^{(m)}(k)$ are the counts in the $k^{th}$ bin of the forecast and mean histograms, normalised to have the same total counts as the observed catalog. Upper and lower quantiles are estimated by comparing the observed statistic
$D_{\text{obs}}=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda_{\text{obs}}^{(m)}(k)+1\right]\right)^{2},$ (14)
with the empirical distribution of $D$ using the repeat simulations, $F_{D}$
$\gamma_{m}=\mathbb{P}(D\leq D_{\text{obs}})=F_{D}(D_{\text{obs}}).$ (15)
Histogram bins of size $\delta_{m}=0.1$ are used across all datasets.
Figure 4: CSEP consistency tests on the ETAS model for the first day ($01/01/2014$) of the testing period in the SCEDC catalog. A total of 10,000 simulations are generated to compute empirical distributions of the test statistics for each of the three consistency tests: (a) Number test, (b) Spatial test, and (c) Magnitude test. The test fails if the observed statistic falls within the rejection region (red), defined by the $0.05$ and $0.95$ quantiles of the distribution.
5.4 Evaluating Multiple Forecasting Periods
Savran et al. (2020) describe how to assess a model’s performance across the multiple days in the testing period. By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator (Gneiting & Katzfuss, 2014). Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast. Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively). The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
Further documentation of how to perform the CSEP evaluation procedure can be found on the platform, where we demonstrate the procedure for the ETAS model. Table 1 reports the benchmark performance scores taken from the quantile plots in Appendix C. The performance of ETAS is higher for the more typical higher magnitude catalogs such as ComCat and SCEDC, whereas it performs worse at the lower magnitude catalogs of QTM_SanJac, QTM_SaltonSea and White. Spatial prediction is consistently the best performing component of the ETAS forecast, whereas earthquake numbers are overpredicted by the model and earthquake magnitudes are generally not well predicted (Figure 9) . All results indicate significant room for improvement beyond the predictive performance of the ETAS model.
Table 1: CSEP consistency tests evaluate the calibration of all daily ETAS forecasts on EarthquakeNPP datasets. A test is performed at the $\alpha=0.05$ significance level on each day in the testing period. The pass rate indicates the success of ETAS across all testing days. By construction quantile scores of the tests should be uniformly distributed if the model is the data generator. The KS-Statistic reports the difference of the quantile distribution to uniform, taken from the quantile plots in Appendix C.
Dataset Number Test Spatial Test Magnitude Test
Pass Rate KS-Statistic Pass Rate KS-Statistic Pass Rate KS-Statistic
ComCat 62.3% 0.392 85.3% 0.128 75.3% 0.318
SCEDC 74.4% 0.161 87.5% 0.123 80.5% 0.153
QTM_SanJac 59.2% 0.461 96.7% 0.145 66.2% 0.406
QTM_SaltonSea 54.2% 0.441 82.1% 0.216 79.0% 0.311
White 17.1% 0.750 98.0% 0.373 25.0% 0.741
6 Discussion and Conclusion
We introduce the EarthquakeNPP datasets to facilitate the benchmarking of NPPs against a community-endorsed ETAS model for earthquake forecasting. These datasets cover various regions of California, representing typical forecasting zones and the data commonly available to forecast issuers. Several datasets use modern methods of detection, which enables the inclusion of much smaller magnitude earthquakes.
In a benchmarking experiment, we compared three NPP models against ETAS and a baseline Poisson process. None of the NPP models outperformed ETAS, indicating that current NPP implementations are not yet suitable for operational earthquake forecasting. ETAS explicitly defines how larger earthquake magnitudes increase the likelihood of future earthquakes in both time and space, with an empirical relationship derived from observational studies (Utsu & Seki, 1955; Utsu, 1970). Since the NPPs lack any direct dependence on magnitudes, this is the likely cause for their performance relative to ETAS. Future implementations should exploit this additional feature for improved temporal and spatial performance. Encouragingly, the comparable temporal performance to ETAS without this additional feature suggests that incorporating magnitude dependence would enhance NPP performance beyond that of ETAS.
EarthquakeNPP supports the earthquake forecast evaluation procedure developed by the Collaboratory for the Study of Earthquake Predictability (CSEP). The procedure replicates how earthquakes forecasts are generated in an operational setting, requiring models to simulate many repeat event sequences over a day-long forecast horizon. Benchmark performance for the ETAS model enables future comparison of NPPs that are implemented for this procedure and enables their promotion to the fully prospective CSEP experiments. Notably, this procedure allows the evaluation of generative NPP models without explicit likelihoods (Yuan et al., 2023; Li et al., ), by assessing their performance over the full trajectory of future events. Probabilistic seismic hazard analysis (PSHA) requires long-term prediction beyond the next-event (Ebrahimian et al., 2014; Gerstenberger et al., 2014), therefore this approach also offers stakeholders a more comprehensive understanding of earthquake hazard than metrics focused on predicting the next event (e.g. RMSE). The procedure also follows the recommendation by Shchur et al. (2021) to move away from next-event point prediction for NPPs.
The EarthquakeNPP datasets, available at https://anonymous.4open.science/r/EarthquakeNPP-2D51, provide a platform for future NPP developments to be benchmarked against these initial results. The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions. Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
Appendix A Earthquake Catalog Data
A.1 Earthquake Catalog Generation
Figure 5: Generating an earthquake catalog involves several key steps: seismic phase picking, magnitude estimation, and the association and location of seismic sources. This process transforms raw waveform data recorded at seismic stations to locations, times, and magnitudes of earthquakes.
Data missingness, referred to in seismology as catalog (in)completeness, is the primary challenge faced with earthquake catalogs. It is an important and unavoidable feature, and is a result of how earthquakes are detected and characterised. Below, we briefly overview the process of generating an earthquake catalog to illustrate the data quality issues. In the subsequent section, we review catalog incompleteness and its potential impact on the performance and evaluation of forecast models.
Seismometers and Seismic Networks. A seismometer is an instrument that detects and records the vibrations caused by seismic waves (Stein & Wysession, 2009; Shearer, 2019). It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities. Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., (Woessner et al., 2010) and references therein). High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
From Waveforms to Phase Picking. The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake. Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes (Allen, 1982). Recent algorithms, such as machine learning classifiers (e.g. Zhu & Beroza, 2019; Lapins et al., 2021) and template-matching (e.g. Ross et al., 2019), can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake. Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event. Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms (Thurber, 1985; Lomax et al., 2000). Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist (e.g. Woessner et al., 2010, and references therein). Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter (1935). The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers. This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons. This led to introduction of the moment magnitude scale (Mw) (Hanks & Kanamori, 1979), which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
$M_{0}=\text{rigidity}\times\text{rupture area}\times\text{slip},$ (16)
where rigidity is a mechanical property of the rock along the fault, rupture area is the area of the fault that slipped, and slip is the distance the fault moved. Mw is determined seismologically via a spectral fitting process to the earthquake waveforms. In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales (e.g. Herrmann & Marzocchi, 2021, and references therein).
A.2 Earthquake Catalog Completeness
All of the EarthquakeNPP datasets are made publicly available by their respective data centers in raw format. However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing. This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ (e.g., Mignan et al., 2011; Mignan & Woessner, 2012).
There are several reasons why an earthquake may not be detected by a seismic network. Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations. Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high (Kagan & Knopoff, 1987; Hainzl, 2022). Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g. when phase arrivals of different events overlap at different stations or the amplitudes of small events are swamped by those of large events. Since catalog incompleteness increases for lower magnitude events, typically the task is to find the value $M_{c}$ above which there is approximately $100\%$ detection probability. Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data. Where NPPs have demonstrated an ability to perform well with incomplete data (Stockman et al., 2023), typically a threshold below the completeness biases classical models such as ETAS (Seif et al., 2017). Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds (e.g. Mancini et al., 2022; Stockman et al., 2023), which we also facilitate in our datasets.
Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law (Gutenberg & Richter, 1936), which states that the distribution of earthquake magnitudes follows an exponential probability density function
$f_{GR}(m)=\beta e^{\beta(m-M_{c})}\ \ :m\geq M_{c}.$ (17)
where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$. Histogram-based approaches, such as the simple Maximum Curvature method (Wiemer & Wyss, 2000) as well as many others (e.g. Herrmann & Marzocchi, 2021, and references therein), identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ (e.g. Schorlemmer & Woessner, 2008). During aftershock sequences, $M_{c}(t)$ can be very high (e.g., Agnew, 2015; Hainzl, 2016b) (See Figure 6a). Thresholding at the maximum value might remove too much data. Instead, modelers either omit particularly incomplete periods during training and testing (Kagan, 1991; Hainzl et al., 2008), model the incompleteness itself (Helmstetter et al., 2006; Werner et al., 2011; Omi et al., 2014; Hainzl, 2016a; b; Mizrahi et al., 2021; Hainzl, 2022), or accept known biases from disregarding this issue (Sornette & Werner, 2005). Spatially, catalogs are less complete farther from the seismic network (Mignan et al., 2011), so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
Figure 6: a) the June 10, 2016 Mw5.2 Borrego Springs earthquake and aftershocks, which occurred on the San Jacinto fault zone and is recorded in the White catalog. An estimate of the magnitude of completeness $M_{c}(t)$ over time using the Maximum Curvature method reveals more incompleteness immediately following the large earthquake. b) magnitude-frequency histograms reveal that truncating the raw White catalog to inside the target region decreases $M_{c}$. Each histogram is fit to the Gutenberg-Richter (GR) law and an estimate of $M_{c}$ for each catalog occurs where the histogram deviates from the (GR) line. c) An estimate of $M_{c}$ for gridded regions of the San Jacinto fault zone, using the raw White catalog.
Appendix B Additional Benchmark Results
Figures 7 and 8 report the temporal and spatial log-likelihood scores of all the benchmarked models on additional datasets. On synthetic data generated by the ETAS model the performance of NPPs mirrors the results on the observational data (Figures 2 and 3). The performance of NPPs is more comparable to ETAS in terms of temporal log-likelihood however they cannot capture the distribution of earthquake locations. Change in temporal performance of models between the ETAS and ETAS_incomplete datasets reveal each model’s robustness to the missing data typically present in earthquake catalogs (See section A.2). Auto-STPP and ETAS reduce in performance upon the removal earthquakes during aftershock sequences, whereas DeepSTPP and NSTPP maintain the same performance indicating a robustness to the data missingness.
On the Japan_Deprecated dataset, whilst ETAS remains the best performing model for spatial prediction, for temporal prediction it performs comparably to NSTPP and is even marginally outperformed by DeepSTPP. This performance can be attributed to the data completeness issues of the Japan_Deprecated dataset (see section 1.1), where the test period is missing all earthquakes bellow magnitude 4.0.
Figure 7: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the additional datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Figure 8: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the additional datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Appendix C CSEP Consistency Tests
Figure 9: Quantile-quantile plots showing the calibration of all daily ETAS forecasts on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White. By construction quantile scores over multiple periods should be uniformly distributed if the model is the data generator. Comparing quantile scores against standard uniform quantiles ($y=x$), highlights discrepancies between the observed data and the forecast. Pass rates of each test are indicated in the legend. The Kolmogorov-Smirnov statistic, quantifies the degree of difference to the uniform distribution.

### v2

@@
-EarthquakeNPP: Benchmark Datasets for Earthquake Forecasting with Neural Point Processes
+EarthquakeNPP: A Benchmark for Earthquake Forecasting with Neural Point Processes
 Abstract
-Classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes for decades.
-Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over classical models.
-However, the currently-used benchmark dataset for NPPs does not represent an up-to-date challenge in the seismological community since it lacks a key earthquake sequence from the region and improperly splits training and testing data.
-Furthermore, initial earthquake forecast benchmarking lacks a comparison to state-of-the-art earthquake forecasting models typically used by the seismological community.
-To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by a credible implementation of the ETAS model.
+For decades, classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes.
+Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over such classical models.
+However, the currently-used benchmark for NPPs does not represent an up-to-date challenge in the seismological community, since it contains data leakage and omits the largest earthquake sequence from the region.
+Additionally, initial earthquake forecasting benchmarks fail to compare NPPs with state-of-the-art forecasting models commonly used in seismology.
+To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by an implementation of the state-of-the-art forecasting model: ETAS.
 The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation.
-In a benchmarking experiment, we compare three spatio-temporal NPPs against ETAS and find that none outperform ETAS in either spatial or temporal log-likelihood.
-These results indicate that current NPP implementations are not yet suitable for practical earthquake forecasting.
-However, EarthquakeNPP will serve as a platform for collaboration between the seismology and machine learning communities with the goal of improving earthquake predictability.
+Benchmarking experiments, using both log-likelihood and generative evaluation metrics widely recognised in seismology, show that none of the five NPPs tested outperform ETAS.
+These findings suggest that current NPP implementations are not yet suitable for practical earthquake forecasting.
+Nonetheless, EarthquakeNPP provides a platform to foster future collaboration between the seismology and machine learning communities.
 1 Introduction
 Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes.
-While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model (Ogata, 1988; 1998), despite significant growth in available data (Takanami et al., 2003; Shelly, 2017; Ross et al., 2019; White et al., 2019; Mousavi et al., 2020; Tan et al., 2021; Mousavi & Beroza, 2023).
-In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019; Jia & Benson, 2019; Chen et al., 2021; Zhou et al., 2022; Zhou & Yu, 2024).
+While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model [53, 54], despite significant growth in available data [79, 72, 62, 90, 51, 80, 50].
+In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility [9, 56, 69, 31, 5, 97, 96].
 While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community.
-The benchmark lacks a key earthquake sequence from the region, fails to recreate an operational setting with proper train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
+The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
 Here, we introduce EarthquakeNPP: a curated collection of datasets designed for benchmarking NPP models in earthquake forecasting, accompanied by a state-of-the-art benchmark model.
@@
 Moreover, employing modern techniques, some datasets include smaller magnitude earthquakes, exploring the potential of numerous small events to enhance forecasting performance through flexible NPPs.
-To unify efforts, we present an operational-level implementation of the ETAS model alongside the datasets, serving as a benchmark for NPPs.
-Although initial benchmarking finds that none of the 3 tested NPP implementations outperform ETAS, EarthquakeNPP aims to serve as a platform for future NPP development.
-The platform facilitates the generative evaluation procedure used for rigorous benchmarking in the seismology community, directing the impact of future NPPs to stakeholders in seismology.
-Access to the dataset collection, along with comprehensive documentation and notebooks, can be found at https://anonymous.4open.science/r/EarthquakeNPP-2D51.
+To unify efforts, we present an operational-level implementation of the ETAS model alongside the datasets, serving as the benchmark for NPPs.
+Although initial benchmarking shows that none of the five tested NPP implementations outperform ETAS, EarthquakeNPP is designed to support ongoing model development and evaluation.
+In addition to the standard log-likelihood metric common in the NPP literature, the platform incorporates the generative evaluation procedures used in seismology for more rigorous benchmarking.
+This ensures that future NPPs (and other models such as time series approaches [88] and Bayesian point processes [68]) can have direct relevance to seismological stakeholders.
+All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP/tree/main.
 1.1 Related Work
-Existing Benchmark Dataset.
+Benchmarking by the NPP Community.
 Chen et al.
-(2021) introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S.
+[5] introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S.
 Geological Survey, focusing on Japan from 1990 to 2020.
 They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset.
-They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake (Mori et al., 2011), the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0.
+They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake [49], the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0.
 This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact.
-Additionally, these events are of significant scientific interest because they provide valuable insights into the earthquake rupture process.
 The dataset segments are divided for training, testing, and validation.
 Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern.
-This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering (Freed, 2005).
-Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards it time.
+This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering [13] - a form of data leakage.
+Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards in time.
 Although earthquakes with magnitudes above 2.5 are considered by Chen et al.
-(2021), following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset.
+[5], following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset.
 For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets.
@@
 Chen et al.
-(2021) benchmark their model against another spatio-temporal model, Neural Jump SDEs (Jia & Benson, 2019), and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark.
-Subsequent papers adopting this benchmark (Zhou et al., 2022; Yuan et al., 2023; Zhou & Yu, 2024) similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
-Temporal-NPP Benchmarking on Earthquake Data.
-Two existing works benchmark NPPs for earthquake forecasting within the seismology community.
-The first by Dascher-Cousineau et al.
-(2023) extends a temporal-only NPP from Shchur et al.
-(2019) to include earthquake magnitudes.
-The second by Stockman et al.
-(2023) extends another temporal-only model by Omi et al.
-(2019a) to target larger magnitude events.
-Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline.
-Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
-Benchmarking within the Seismology Community.
-Model comparison has been crucial in the development of earthquake forecasting models since their inception (Kagan & Knopoff, 1987; Ogata, 1988).
-The Collaboratory for the Study of Earthquake Predictability (CSEP) (Michael & Werner, 2018; Schorlemmer et al., 2018; Savran et al., 2022; Iturrieta et al., 2024) (https://cseptesting.org/ ) aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally.
+[5] benchmark their model against another spatio-temporal model, Neural Jump SDEs [31], and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark.
+Subsequent papers adopting this benchmark [97, 94, 96] similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
+Benchmarking by the Seismology Community.
+Model comparison has been crucial in the development of earthquake forecasting models since their inception [34, 53].
+The Collaboratory for the Study of Earthquake Predictability (CSEP) [41, 66, 64, 30] [https://cseptesting.org/ ] aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally.
 CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day).
@@
 The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
-1.2 Scope of this work
-Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development (Ogata, 1988; Harte, 2015).
-Our platform uses this metric in the NPP benchmarking experiment and provides detailed guidance on CSEP’s simulation-based procedure, enabling future NPP implementations and evaluations within CSEP experiments.
-The goal of this work is to allow Machine Learning researchers to have seismological impact by defining a baseline target for which NPP models can be compared to state-of-the-art domain-based models.
-NPPs that can generate log-likelihoods are in scope, whilst those that do not (e.g.
-Yuan et al., 2023; Li et al., 2023) are out of scope because a valid score does not currently exist.
-The popular next-event point prediction metrics (e.g.
-Root Mean Square Error (RMSE) and related scores) are considered to be flawed and misleading for seismological prediction (Hodson, 2022), because the predictive distribution is strongly skewed and therefore far from Gaussian.
-To have seismological relevance, authors of NPP models are challenged to implement long-term predictions using CSEP’s evaluation procedure, benchmarking against the reported performance for the ETAS model.
+Two existing works benchmark NPPs for earthquake forecasting within the seismology community.
+The first by Dascher-Cousineau et al.
+[8] extends a temporal-only NPP from Shchur et al.
+[69] to include earthquake magnitudes.
+The second by Stockman et al.
+[77] extends another temporal-only model by Omi et al.
+[56] to target larger magnitude events.
+Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline.
+Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
 2 Background
 2.1 Spatio-Temporal Point Processes
-A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\in\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$.
+A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$.
 This process is typically defined by a non-negative conditional intensity function
-$\lambda(t,\mathbf{x}|\mathcal{H}_{t}):=\lim_{\Delta t,\Delta\mathbf{x}\rightarrow 0}\frac{\mathbb{E}\left[N([t,t+\Delta t)\times B(\mathbf{x},\Delta\mathbf{x})|\mathcal{H}_{t}\right]}{|B(\mathbf{x},\Delta\mathbf{x})|},$ (1)
+$\lambda(t,\mathbf{x}|\mathcal{H}_{t}):=\lim_{\Delta t,\Delta\mathbf{x}\rightarrow 0}\frac{\mathbb{E}\left[N([t,t+\Delta t)\times B(\mathbf{x},\Delta\mathbf{x})|\mathcal{H}_{t}\right]}{\Delta t|B(\mathbf{x},\Delta\mathbf{x})|},$ (1)
 where $\mathcal{H}_{t}=\{(t_{i},\mathbf{x}_{i})|t_{i}<t\}$ denotes the history of events preceding time $t$ and $|B(\mathbf{x},\Delta\mathbf{x})|$ is the Lebesgue measure of the ball $B(\mathbf{x},\Delta\mathbf{x})$ with radius $\Delta\mathbf{x}$.
 Given we observe a history of events up to $t_{i}$, the probability density function (pdf) of observing an event at time $t$ and location $\mathbf{x}$ is given by
-[…]
-Most models specify the conditional intensity function, though some (e.g.
-Shchur et al., 2019; Chen et al., 2021; Yuan et al., 2023), directly model this pdf.
-Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $[T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
-[…]
+$p(t,\mathbf{x}|\mathcal{H}_{t_{i}})=\lambda(t,\mathbf{x}|\mathcal{H}_{t_{i}})\cdot\exp\left(-\int_{t_{i}}^{t}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds\right).$ (2)
+Most models specify the conditional intensity function, though some [e.g.
+69, 5, 94] directly model this pdf.
+Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
+$\log p(\mathcal{H}_{T})=\underbrace{\sum_{i=0}^{n}\log\lambda(t_{i}|\mathcal{H}_{t_{i}})-\int_{T_{0}}^{T_{1}}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds}_{\text{Temporal log-likelihood}}+\underbrace{\sum_{i=0}^{n}\log f(\mathbf{x}_{i}|t_{i},\mathcal{H}_{t_{i}})}_{\text{Spatial log-likelihood}},$ (3)
 where the decomposition of the spatio-temporal conditional intensity function, $\lambda(t_{i},\mathbf{x}_{i}|\mathcal{H}_{t_{i}})=\lambda(t_{i}|\mathcal{H}_{t_{i}})\cdot f(\mathbf{x}_{i}|t_{i},\mathcal{H}_{t_{i}})$, allows the log-likelihood to be written as contributions from the temporal and spatial components.
@@
 For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric.
-This is consistent with a wealth of literature in the seismology community (see Zechar et al., 2010, and references therein) as well as the wider general point process literature (Daley & Vere-Jones, 2004), which now more recently includes neural point processes (Shchur et al., 2021).
+This is consistent with a wealth of literature in the seismology community [see 95, and references therein] as well as the wider general point process literature [7], which now includes neural point processes [70].
 The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident.
-Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $[T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function.
-Comparing the mean log-likelihood per event provides the information gain from one model to another (Daley & Vere-Jones, 2004).
+Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function.
+Comparing the difference in mean log-likelihood per event provides the information gain from one model to another [7].
+Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting [47] and established benchmarking experiments (CSEP) [81, 60].
+The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time [33].
+Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin — a delay that could miss critical, potentially damaging events.
 2.2 ETAS
-The Epidemic Type Aftershock Sequence (ETAS) model (Ogata, 1998) is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space.
-It has been adopted for operational earthquake forecasting by government agencies in California (Milner et al., 2020), New-Zealand (Christophersen et al., 2017), Italy (Spassiani et al., 2023), Japan (Omi et al., 2019b) and Switzerland (Mizrahi et al., 2024), and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments (e.g.
-Woessner et al., 2011; Rhoades et al., 2018; Taroni et al., 2018; Cattania et al., 2018; Mancini et al., 2019; 2020; 2022).
+The Epidemic Type Aftershock Sequence (ETAS) model [54] is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space.
+It has been adopted for operational earthquake forecasting by government agencies in California [44], New-Zealand [6], Italy [74], Japan [57] and Switzerland [48], and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments [e.g.
+93, 60, 81, 4, 38, 39, 40].
 The general formulation of the model is
-$\lambda(t,\mathbf{x}|\mathcal{H}_{t};\mathbf{\theta})=\mu+\sum_{i:t_{i}<t}g(t-t_{i},||\mathbf{x}-\mathbf{x}_{i}||^{2}_{2},m_{i}),$ (4)
+[…]
 where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event.
@@
 This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g.
-via dynamic wave triggering (Brodsky & van der Elst, 2014) and static stress triggering (Gomberg, 2018; Mancini et al., 2020).
-[…]
+via dynamic wave triggering [3] and static stress triggering [16, 39].
+Although ETAS can be fit by maximizing the log-likelihood function directly, parameter estimation is typically performed by simultaneously estimating the branching structure $\mathbf{B}$.
+Veen and Schoenberg [87] developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
 $\mathbf{\theta}^{(k+1)}=\operatorname*{arg\,max}_{\mathbf{\theta}}\mathbb{E}_{\mathbf{B}\sim p(\cdot|\mathcal{H}_{T_{1}},\mathbf{\theta}^{(k)})}\left[\log p(\mathcal{H}_{T_{1}},\mathbf{B}|\mathbf{\theta})\right].$ (5)
-This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation and simultaneously estimates the causal structure.
-The formulation of the ETAS model we present with the EarthquakeNPP datasets is implemented in the etas python package by Mizrahi et al. (2022). It defines the triggering kernel as
-$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{cut})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{cut})}\right)^{1+\rho}},$ (6)
+This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation, and simultaneously distinguishes background events from triggered events.
+Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP.
+Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans.
+In red is a fault map from the GEM Global Active Faults Database [78].
+The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. [46]. It defines the triggering kernel as
+$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{c})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{c})}\right)^{1+\rho}},$ (6)
 where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$.
+This triggering kernel is derived from statistical distributions found through decades of observational studies [84, 83, 85] and several of the learnable parameters have been linked to physical properties of the earthquake rupture process [85, 29].
 3 EarthquakeNPP Datasets
-The EarthquakeNPP datasets encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021.
-California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms (Gerstenberger et al., 2004; Field, 2007; Field et al., 2021).
-It encompasses the San Andreas fault plate boundary system (Zoback et al., 1987) and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
-Notebooks to access and preprocess these public datasets along with the associated benchmarking experiment are publicly accessible at https://anonymous.4open.science/r/EarthquakeNPP-2D51, accompanied by more detailed documentation for each dataset.
+The EarthquakeNPP datasets (Figure 1) encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021.
+California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms [14, 11, 12].
+It encompasses the San Andreas fault plate boundary system [99] and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
+Notebooks to access and preprocess these public datasets along with the associated benchmarking experiment are publicly accessible at https://github.com/ss15859/EarthquakeNPP/tree/main, accompanied by more detailed documentation for each dataset.
 A summary of how earthquake datasets are generated, along with the associated challenges of using earthquake catalog data can be found in Appendix A.
-The following subsections provide a short overview of each EarthquakeNPP dataset.
-3.1 ANSS Comprehensive Earthquake Catalog (ComCat)
-The U.S.
+Table 1 provides a short summary of each EarthquakeNPP dataset.
+Table 1: Summary of EarthquakeNPP datasets, including: region, dataset development, magnitude threshold ($\mathbf{M}_{\mathbf{c}}$), number of training (combined with validation) events, and number of testing events.
+The chronological partitioning of training, validation, and testing periods is also detailed.
+An auxiliary (burn-in) period begins from the Start date, followed by the respective starts of the training, validation, and testing periods.
+All dates are given as 00:$0$ UTC on January $1^{\text{st}}$, unless noted (* refers to 00:$0$ UTC on January $17^{\text{th}}$).
+Finally, we give our purpose for including each dataset.
+ComCat SCEDC White QTM
+Region Whole of California Southern California San Jacinto Fault-Zone QTM_SanJac:
+San Jacinto Fault-Zone,
+QTM_SaltonSea: Salton Sea
+Development The U.S.
 Geological Survey (USGS) National Earthquake Information Center (NEIC) monitors global earthquakes (Mw 4.5 or larger) and provides complete seismic monitoring of the United States for all significant earthquakes (> Mw 3.0 or felt).
 Its contributing seismic networks have produced the Advanced National Seismic System (ANSS) Comprehensive Catalog of Earthquake Events and Products.
-We focus on the California region defined by Schorlemmer & Gerstenberger (2007), with a test period consistent with CSEP experiments (Zechar et al., 2013).
-3.2 Southern California Earthquake Data Center (SCEDC) Catalog
-The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California (Hutton et al., 2010) since the Caltech Seismological Laboratory began routine operations in 1932.
+The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California [28] since the Caltech Seismological Laboratory began routine operations in 1932.
 Significant network improvements since the 1970s and 1980s reduced the catalog completeness from Mw 3.25 to Mw 1.8.
-We use three magnitude thresholds (Mw 2.0, 2.5, 3.0) to explore the effect of truncation on forecasting model performance.
-Training includes the Mw 7.3 Landers and 1999 Mw 7.1 Hector Mine earthquakes, while testing involves the 2019 Mw 7.1 Ridgecrest sequence.The USGS utilizes both ComCat and SCEDC datasets in the aftershock forecasts they release to the public.
-The inclusion of these datasets determines whether NPPs can exploit the datasets currently being used for operational forecasting.
-3.3 Detailed Earthquake Catalog for the San Jacinto Fault-Zone Region
 White et al.
-(2019) created an enhanced catalog focusing on the San Jacinto fault region, using an dense seismic network in Southern California.
-This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion (White et al., 2019).
-The training window includes the 2010 Mw 5.4 Borrego Springs and 2013 ML 4.7 Anza Borrego earthquakes.
-This catalog is named White after the authors.
-Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP.
-Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans.
-In red is a fault map from the GEM Global Active Faults Database (Styron & Pagani, 2020).
-3.4 Quake Template Matching (QTM) Catalog
+[90] created an enhanced catalog focusing on the San Jacinto fault region, using a dense seismic network in Southern California.
+This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion [90].
 Using data collected by the SCSN, Ross et al.
-(2019) generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
-The catalog contains 1.81 million earthquakes complete down to Mw 0.3.
-Following Dascher-Cousineau et al.
-(2023), we use a more conservative completeness estimate of Mw 1.0 and split the catalog into two focus regions: the San Jacinto fault region and the Salton Sea.
-The inclusion of White, QTM_SanJac and QTM_SaltonSea determines whether very low magnitude earthquakes exhibit different behaviour to those of the larger earthquakes, an assumption which ETAS makes.
-3.5 Additional Datasets
-Beyond the official EarthquakeNPP datasets, we include 3 further datasets that either provide additional scientific insight or continuity from previous benchmarking works.
-[…]
-$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (7)
-where $M$ is the mainshock magnitude.
-Events below this threshold are removed using mainshocks of Mw 5.2 and above.
-The inclusion of these datasets allows us to test whether NPPs are inhibited by data missingness to the same extent that ETAS is.
-Deprecated Catalog of Japan.
-To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al.
-(2021), however with a chronological train-test split and without removing any supposed outlier events.
-To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
+[62] generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
+$\mathbf{M}_{\mathbf{c}}$ Mw 2.5 SCEDC_20: Mw 2.0,
+SCEDC_25: Mw 2.5,
+SCEDC_30: Mw 3.0
+Mw 0.6 Mw 1.0
+# Train/Test
+Events
+79,037 / 23,059 SCEDC_20:
+128,265 / 14,351,
+SCEDC_25:
+43,221 / 5,466,
+SCEDC_30:
+12,426 / 2,065
+38,556 / 26,914 QTM_SanJac:
+18,664 / 4,837,
+QTM_SanJac:
+44,042 / 4,393
+Start-Train-Val-Test-End 1971-1981-1998-2007-2020 1981-1985-2005-2014-2020 2008-2009-2014-2017-2021 2008-2009-2014-2016-2018
+Purpose Example of data currently in use for operational
+forecasting (USGS utilizes ComCat in aftershock forecasts they release to the public.)
+Three magnitude thresholds (Mw 2.0, 2.5, 3.0) explore the effect of truncation on forecasting model performance.
+To explore if newly detected low magnitude earthquakes contain additional predictive information.
+To explore if newly detected low magnitude earthquakes contain additional predictive information (with different detection methodology).
 4 Benchmarking Experiment
-We now use EarthquakeNPP to benchmark three spatio-temporal NPPs with prior positive claims on earthquake forecasting.
-Neural Spatio-Temporal Point Process (NSTPP) (Chen et al., 2021): a pdf based NPP that parameterizes the spatial pdf with continuous-time normalizing flows (CNFs).
-We use their Attentive CNF model for its computational efficiency and overall performance versus their other model Jump CNF (Chen et al., 2021).
-Deep Spatio-Temporal Point Process (Deep-STPP) (Zhou et al., 2022): a conditional intensity function based NPP that constructs a non parametric space-time intensity function governed by a deep latent process.
-The intensity function enjoys a closed form integration, avoiding the need for numerical approximation.
-Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) (Zhou & Yu, 2024): a conditional intensity function based NPP which jointly models the 3D space-time integral of the intensity along with its derivative (the intensity function) using a dual network approach.
-The benchmark is against the ETAS model defined in section 2.2, as well as a homogeneous Poisson process.
-The Poisson model is fit to events in the auxiliary, training and validation windows to provide a baseline score against which to compare all four other models.
-Validation is typically not part of the estimation procedure for ETAS, so it is fit using the combined training and validation windows.
-NPPs follow the standard training/validation/testing procedure of machine learning.
-When possible, a model’s likelihood for training, validation, and testing can depend on events occurring before the splits through memory in its history.
-The exception is NSTPP, lacking a direct dependency on prior events.
-Nonetheless, its likelihood is evaluated on the same events as the other models.
+We use EarthquakeNPP to benchmark five spatio-temporal Neural Point Processes (NPPs), each with prior positive claims in earthquake forecasting, against the ETAS model described in Section 2.2.
+Neural Spatio-Temporal Point Process (NSTPP) [5]: A probability density function (pdf)-based NPP that parametrizes the spatial pdf with continuous-time normalizing flows (CNFs).
+We evaluate their Attentive CNF model due to its superior computational efficiency and overall performance compared to the Jump CNF model [5].
+Deep Spatio-Temporal Point Process (DeepSTPP) [97]: A conditional intensity function-based NPP that constructs a non-parametric space-time intensity function driven by a deep latent process.
+This model features a closed-form intensity function, eliminating the need for numerical approximations.
+Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) [96]: A conditional intensity function-based NPP that jointly models the 3D space-time integral of the intensity and its derivative (the intensity function) using a dual network approach.
+Spatio-temporal Diffusion Point Process (DSTPP) [94]: A generative NPP that does not have a likelihood function. DSTPP employs diffusion models to capture complex spatio-temporal dynamics.
+Score Matching-based Pseudolikelihood Estimation of Neural Marked Spatio-Temporal Point Process (SMASH)[36]: A generative NPP that also lacks a likelihood function.
+SMASH adopts a normalization-free objective by estimating the pseudolikelihood of marked STPPs through score-matching.
+4.1 Likelihood Evaluation
+Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development [53, 23].
+Other traditional next-event metrics like Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are misleading for earthquake forecasting [27], as earthquake occurrence follows power law distributions [33, 10] that are heavy-tailed, making the errors non-Gaussian and non-Laplacian — contrary to the assumptions underlying RMSE and MAE (see Section G).
+For the three models with valid likelihood functions (NSTPP, DeepSTPP, and AutoSTPP), we present the mean log-likelihood scores in Figures 2 and 3.
+These scores are compared alongside the ETAS model (Section 2.2) and a homogeneous Poisson process.
+The Poisson model is fit to events in the auxiliary, training, and validation windows to provide a baseline score for comparison.
+Unlike the NPPs, which follow the standard machine learning procedure of training, validation, and testing, ETAS does not typically incorporate validation in its estimation procedure.
+Thus, it is fit using both the training and validation windows combined.
+For NPPs, the likelihood for training, validation, and testing depends on events occurring before the respective splits through memory in their history.
+The exception is NSTPP, which lacks a direct dependency on prior events.
+Nevertheless, its likelihood is evaluated on the same data as the other models.
 The definition of the ETAS model (equation 4) specifies how the magnitudes of earthquakes in the history contribute towards the intensity function.
 This earthquake magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modeling choices beyond the scope of this work.
-Figures 2 and 3 report the temporal and spatial log-likelihood scores of all models on the EarthquakeNPP datasets.
-The ETAS model achieves the highest temporal and spatial log-likelihood across all datasets, with some NPP models achieving comparable temporal performance on ComCat, QTM_SaltonSea, QTM_SanJac, and White catalogs.
-Amongst the NPP models, Deep-STPP generally performs best in terms of temporal log-likelihood, whereas AutoSTPP performs best in terms of spatial log-likelihood.
-The improved relative temporal performance of all NPPs compared to ETAS as the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, as well as the comparable performance to ETAS using the low magnitude catalogs QTM_SaltonSea, QTM_SanJac, and White, indicates that low magnitude earthquakes provide valuable predictive information for NPPs.
-Appendix B provides further results using the additional datasets.
+The ETAS model consistently achieves the highest temporal log-likelihood, with NPPs performing comparably or, in some cases, marginally better, except at the higher magnitude thresholds of the SCEDC catalog.
+Among the NPPs, AutoSTPP and NSTPP perform well across several datasets, though their performance is more variable than that of DeepSTPP, which demonstrates consistent, comparable performance to ETAS.
+Closer examination of model performance over time (Figure 9) reveals that relative performance to ETAS is poorest during large earthquake sequences.
+This is likely due to ETAS leveraging the magnitude feature of the data, which enables it to handle these sequences effectively.
+Conversely, model performance is strongest during "background" periods, when no large earthquakes occur.
+During these periods, ETAS models the background with a constant rate, while the NPPs improve upon this by capturing the non-stationary nature of the background data.
+This effect is most pronounced in the ComCat dataset, which includes more complex physical processes, such as those near the Mendocino Triple Junction [24].
+The improved relative temporal performance of all NPPs compared to ETAS, particularly when the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, indicates that low magnitude earthquakes provide valuable predictive information for NPPs.
+ETAS consistently outperforms all NPPs in spatial log-likelihood.
+As with temporal performance, relative performance to ETAS is weakest during large earthquake periods, likely due to the absence of a magnitude feature in the NPPs.
+AutoSTPP achieves the highest spatial log-likelihood, attributed to its ability to capture anisotropic Hawkes kernels (see Figure 2 of Zhou and Yu [96]), which are commonly observed in earthquake data [58].
 Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
@@
 Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
-5 CSEP Consistency Tests
+4.2 CSEP Consistency Tests
 EarthquakeNPP supports the earthquake forecast evaluation protocol developed by the Collaboratory for the Study of Earthquake Predictability (CSEP).
-In this procedure a model generates 24-hour forecasts through $10,000$ repeat simulations of earthquake sequences at the beginning of every day in the testing period.
-This procedure exactly mimics how earthquake forecasts are generated in an operational setting (van der Elst et al., 2022).
-Models can then be evaluated by comparing the observed sequence with the distribution over model simulations.
-Three test statistics target the temporal, spatial and magnitude components of the forecasts, where a test is failed if the observed statistic falls within a pre-defined rejection region (Figure 4).
-We demonstrate this procedure for the ETAS model and report performance scores as a benchmark for future immplementations of NPPs.
-5.1 Number Test
-The number test evaluates the temporal component of the forecast by checking the consistency of the forecasted number of events, $N$ with those observed in the forecast horizon, $N_{\text{obs}}$.
-Upper and lower quantiles are estimated using the empirical cumulative distribution from the repeat simulations, $F_{N}$,
-$\delta_{1}$ $=\mathbb{P}(N\geq N_{\text{obs}})=1-F_{N}(N_{\text{obs}}-1)$ (8)
-$\delta_{2}$ $=\mathbb{P}(N\leq N_{\text{obs}})=F_{N}(N_{\text{obs}}).$ (9)
-5.2 Spatial Test
-To evaluate the spatial component of the forecast, a test statistic aggregates the forecasted rates of earthquakes over a regular grid,
-$S=\left[\sum_{i=1}^{N}\log\hat{\lambda}(k_{i})\right]N^{-1},$ (10)
-where $\hat{\lambda}(k_{i})$ is the approximate rate in the cell $k$ where the $i^{th}$ event is located. Upper and lower quantiles are estimated by comparing the observed statistic
-$S_{\text{obs}}=\left[\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}(k_{i})\right]N_{\text{obs}}^{-1},$ (11)
-with the empirical cumulative distribution of $S$ using the repeat simulations, $F_{S}$
-$\gamma_{s}=\mathbb{P}(S\leq S_{\text{obs}})=F_{S}(S_{\text{obs}}).$ (12)
-The grid is constructed from $\{0.1\degree,0.05\degree,0.01\degree\}$ squares for ComCat, SCEDC and {QTM_Salton_Sea, QTM_SanJac, White} respectively.
-5.3 Magnitude Test
-To evaluate the earthquake magnitude component of the forecast, a test statistic compares the histogram of a forecast’s magnitudes $\Lambda^{(m)}$, against the mean histogram over all forecasts $\bar{\Lambda}^{(m)}$,
-$D=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda^{(m)}(k)+1\right]\right)^{2},$ (13)
-where $\Lambda^{(m)}(k)$ and $\bar{\Lambda}^{(m)}(k)$ are the counts in the $k^{th}$ bin of the forecast and mean histograms, normalised to have the same total counts as the observed catalog.
-Upper and lower quantiles are estimated by comparing the observed statistic
-$D_{\text{obs}}=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda_{\text{obs}}^{(m)}(k)+1\right]\right)^{2},$ (14)
-with the empirical distribution of $D$ using the repeat simulations, $F_{D}$
-$\gamma_{m}=\mathbb{P}(D\leq D_{\text{obs}})=F_{D}(D_{\text{obs}}).$ (15)
-Histogram bins of size $\delta_{m}=0.1$ are used across all datasets.
-Figure 4: CSEP consistency tests on the ETAS model for the first day ($01/01/2014$) of the testing period in the SCEDC catalog.
-A total of 10,000 simulations are generated to compute empirical distributions of the test statistics for each of the three consistency tests: (a) Number test, (b) Spatial test, and (c) Magnitude test.
-The test fails if the observed statistic falls within the rejection region (red), defined by the $0.05$ and $0.95$ quantiles of the distribution.
-5.4 Evaluating Multiple Forecasting Periods
-Savran et al.
-(2020) describe how to assess a model’s performance across the multiple days in the testing period.
-By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator (Gneiting & Katzfuss, 2014).
-Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast.
-Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively).
-The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
-Further documentation of how to perform the CSEP evaluation procedure can be found on the platform, where we demonstrate the procedure for the ETAS model.
-Table 1 reports the benchmark performance scores taken from the quantile plots in Appendix C.
-The performance of ETAS is higher for the more typical higher magnitude catalogs such as ComCat and SCEDC, whereas it performs worse at the lower magnitude catalogs of QTM_SanJac, QTM_SaltonSea and White.
-Spatial prediction is consistently the best performing component of the ETAS forecast, whereas earthquake numbers are overpredicted by the model and earthquake magnitudes are generally not well predicted (Figure 9) .
-All results indicate significant room for improvement beyond the predictive performance of the ETAS model.
-Table 1: CSEP consistency tests evaluate the calibration of all daily ETAS forecasts on EarthquakeNPP datasets.
+In this procedure, a model generates 24-hour forecasts through 10,000 repeated simulations of earthquake sequences at the beginning of each day in the testing period.
+This approach mirrors how earthquake forecasts are produced in operational settings [86].
+Models are then evaluated by comparing the observed sequence with the distribution of forecasts generated by the simulations.
+Four test statistics assess the temporal, spatial, likelihood, and magnitude components of the forecasts.
+A test is considered failed if the observed statistic falls within a pre-defined rejection region (Figure 11).
+We apply this evaluation procedure to the two generative NPPs (DSTPP and SMASH) alongside ETAS (Table 2) and present a case study on the 2010 M7.2 El Mayor-Cucapah earthquake, using the forecasts from these models (Figure 4).
+Appendix E provides an introduction to the CSEP consistency tests, with further details found at https://cseptesting.org/.
+Table 2: CSEP consistency tests evaluate the calibration of daily forecasts from three models (ETAS, SMASH, DSTPP) on EarthquakeNPP datasets.
 A test is performed at the $\alpha=0.05$ significance level on each day in the testing period.
-The pass rate indicates the success of ETAS across all testing days.
-By construction quantile scores of the tests should be uniformly distributed if the model is the data generator.
-The KS-Statistic reports the difference of the quantile distribution to uniform, taken from the quantile plots in Appendix C.
-Dataset Number Test Spatial Test Magnitude Test
-Pass Rate KS-Statistic Pass Rate KS-Statistic Pass Rate KS-Statistic
-ComCat 62.3% 0.392 85.3% 0.128 75.3% 0.318
-SCEDC 74.4% 0.161 87.5% 0.123 80.5% 0.153
-QTM_SanJac 59.2% 0.461 96.7% 0.145 66.2% 0.406
-QTM_SaltonSea 54.2% 0.441 82.1% 0.216 79.0% 0.311
-White 17.1% 0.750 98.0% 0.373 25.0% 0.741
+The pass rate indicates the proportion of testing days with non-rejected hypotheses.
+If the model is the data generator, quantile scores should be uniformly distributed.
+The KS-Statistic quantifies deviation from uniformity (see Appendix E).
+ETAS is the only model that forecasts earthquake magnitudes, so is the only model evaluated with the magnitude test.
+Dataset Model Number Test Spatial Test Pseudo Likelihood Test Magnitude Test
+Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat.
+ComCat ETAS 85.1% 0.222 92.0% 0.029 97.6% 0.128 93.8% 0.113
+SMASH 72.4% 0.212 68.6% 0.217 87.6% 0.134
+DSTPP 70.4% 0.116 88.6% 0.070 86.3% 0.138
+SCEDC ETAS 81.7% 0.347 88.3% 0.119 95.9% 0.233 90.0% 0.256
+SMASH 59.9% 0.602 51.1% 0.471 68.0% 0.611
+DSTPP 0.0% 0.840 6.1% 0.935 0.8% 0.988
+QTM_SanJac ETAS 88.8% 0.151 91.7% 0.095 96.6% 0.123 94.8% 0.076
+SMASH 67.8% 0.290 55.6% 0.385 53.4% 0.584
+DSTPP 78.4% 0.110 85.7% 0.240 85.3% 0.136
+QTM_SaltonSea ETAS 77.8% 0.210 90.9% 0.206 96.4% 0.119 90.6% 0.136
+SMASH 58.6% 0.486 53.6% 0.371 73.7% 0.451
+DSTPP 69.6% 0.154 88.8% 0.136 85.6% 0.127
+White ETAS 83.4% 0.167 86.6% 0.225 90.8% 0.233 88.8% 0.052
+SMASH 61.3% 0.274 84.5% 0.150 67.7% 0.246
+DSTPP 0.0% 0.987 30.9% 0.691 32.3% 0.892
+ETAS consistently performs best across all datasets and tests.
+It achieves the highest pass rates and lowest KS statistics, indicating strong calibration and reliability.
+SMASH shows moderate performance, often outperforming DSTPP but trailing ETAS.
+Its results vary more across datasets and tests, with occasional strengths (e.g.
+in White for spatial KS).
+DSTPP generally performs worse, with lower pass rates and higher KS statistics, especially for the SCEDC and White datasets.
+However, it achieves relatively good spatial calibration in some cases (e.g., ComCat).
+Figure 4: Forecasts from ETAS, SMASH, and DSTPP during the 2010 M7.2 El Mayor-Cucapah earthquake contained in the ComCat dataset.
+Top: Spatial forecasts for the day following the mainshock.
+ETAS accurately captures the primary aftershock zone along the Laguna Salada fault system.
+SMASH produces smoother forecasts with broader spatial spread, while DSTPP concentrates its probability mass north of the mainshock epicenter.
+Bottom: Cumulative earthquake counts over time, with magnitudes shown as scaled orange circles.
+Forecast number distributions from each model are plotted with 95% confidence intervals.
+All models initially underestimate aftershock activity.
+ETAS and SMASH begin to recover after the first week, whereas DSTPP continues to systematically underpredict event counts throughout the sequence.
 5 Discussion and Conclusion
-We introduce the EarthquakeNPP datasets to facilitate the benchmarking of NPPs against a community-endorsed ETAS model for earthquake forecasting.
-These datasets cover various regions of California, representing typical forecasting zones and the data commonly available to forecast issuers.
-Several datasets use modern methods of detection, which enables the inclusion of much smaller magnitude earthquakes.
-In a benchmarking experiment, we compared three NPP models against ETAS and a baseline Poisson process.
-None of the NPP models outperformed ETAS, indicating that current NPP implementations are not yet suitable for operational earthquake forecasting.
-ETAS explicitly defines how larger earthquake magnitudes increase the likelihood of future earthquakes in both time and space, with an empirical relationship derived from observational studies (Utsu & Seki, 1955; Utsu, 1970).
-Since the NPPs lack any direct dependence on magnitudes, this is the likely cause for their performance relative to ETAS.
-Future implementations should exploit this additional feature for improved temporal and spatial performance.
-Encouragingly, the comparable temporal performance to ETAS without this additional feature suggests that incorporating magnitude dependence would enhance NPP performance beyond that of ETAS.
-EarthquakeNPP supports the earthquake forecast evaluation procedure developed by the Collaboratory for the Study of Earthquake Predictability (CSEP).
-The procedure replicates how earthquakes forecasts are generated in an operational setting, requiring models to simulate many repeat event sequences over a day-long forecast horizon.
-Benchmark performance for the ETAS model enables future comparison of NPPs that are implemented for this procedure and enables their promotion to the fully prospective CSEP experiments.
-Notably, this procedure allows the evaluation of generative NPP models without explicit likelihoods (Yuan et al., 2023; Li et al., ), by assessing their performance over the full trajectory of future events.
-Probabilistic seismic hazard analysis (PSHA) requires long-term prediction beyond the next-event (Ebrahimian et al., 2014; Gerstenberger et al., 2014), therefore this approach also offers stakeholders a more comprehensive understanding of earthquake hazard than metrics focused on predicting the next event (e.g.
-RMSE).
-The procedure also follows the recommendation by Shchur et al.
-(2021) to move away from next-event point prediction for NPPs.
-The EarthquakeNPP datasets, available at https://anonymous.4open.science/r/EarthquakeNPP-2D51, provide a platform for future NPP developments to be benchmarked against these initial results.
-The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions.
+We introduce EarthquakeNPP, a benchmarking platform designed to evaluate Neural Point Process (NPP) models against the state-of-the-art ETAS model for earthquake forecasting.
+The platform hosts datasets from diverse regions of California, both standard forecasting zones and datasets that incorporate modern detection techniques.
+We establish two evaluation frameworks tailored to seismology: standard log-likelihood metrics and the generative consistency tests developed by the Collaboratory for the Study of Earthquake Predictability (CSEP), ensuring that successful models can be directly relevant to operational forecasting.
+In benchmarking five NPP models against ETAS, we found that none outperformed the baseline, suggesting current NPPs are not yet suitable for operational use.
+While NPPs often perform well during low-activity background periods, they struggle during active phases following large earthquakes.
+This may be due to their lack of explicit magnitude dependence — a key feature in ETAS — and their limited memory, as NPPs encode only a fixed-length history.
+In contrast, ETAS sums over all past events, allowing long-past earthquakes to influence future rates.
+Future NPPs should incorporate magnitude information and develop mechanisms for longer-term memory.
+EarthquakeNPP, available at https://github.com/ss15859/EarthquakeNPP/tree/main, provides a platform for future NPP developments to be benchmarked against these initial results.
+The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions.
 Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
+Acknowledgments and Disclosure of Funding
+This project is funded by Compass - Centre for Doctoral Training in Computational Statistics and Data Science (EPSRC Grant Ref EP/S023569/1).
+Compass is funded by United Kingdom Research and Innovation (UKRI) through the Engineering and Physical Sciences Research Council (EPSRC), https://www.ukri.org/councils/epsrc.
+This project also has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant 821115, Real-time earthquake rIsk reduction for a reSilient Europe (RISE), http://www.rise-eu.org) and by United States Geological Survey (USGS) EHP grant G24AP00059.
 Appendix A Earthquake Catalog Data
@@
 Seismometers and Seismic Networks.
-A seismometer is an instrument that detects and records the vibrations caused by seismic waves (Stein & Wysession, 2009; Shearer, 2019).
+A seismometer is an instrument that detects and records the vibrations caused by seismic waves [75, 71].
 It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities.
-Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., (Woessner et al., 2010) and references therein).
+Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., [92] and references therein).
 High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
@@
 The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake.
-Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes (Allen, 1982).
-Recent algorithms, such as machine learning classifiers (e.g.
-Zhu & Beroza, 2019; Lapins et al., 2021) and template-matching (e.g.
-Ross et al., 2019), can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
+Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes [2].
+Recent algorithms, such as machine learning classifiers [e.g.
+98, 35] and template-matching [e.g.
+62], can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
 Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake.
 Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event.
-Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms (Thurber, 1985; Lomax et al., 2000).
-Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist (e.g.
-Woessner et al., 2010, and references therein).
+Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms [82, 37].
+Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist [e.g.
+92, and references therein].
 Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
-Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter (1935).
+Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter [61].
 The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers.
 This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons.
-This led to introduction of the moment magnitude scale (Mw) (Hanks & Kanamori, 1979), which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
+This led to introduction of the moment magnitude scale (Mw) [22], which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
 $M_{0}=\text{rigidity}\times\text{rupture area}\times\text{slip},$ (7)
@@
 Mw is determined seismologically via a spectral fitting process to the earthquake waveforms.
-In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales (e.g.
-Herrmann & Marzocchi, 2021, and references therein).
+In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales [e.g.
+26, and references therein].
 A.2 Earthquake Catalog Completeness
@@
 However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing.
-This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ (e.g., Mignan et al., 2011; Mignan & Woessner, 2012).
+This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ [e.g., 42, 43].
 There are several reasons why an earthquake may not be detected by a seismic network.
 Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations.
-Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high (Kagan & Knopoff, 1987; Hainzl, 2022).
+Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high [34, 20].
 Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g.
@@
 Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data.
-Where NPPs have demonstrated an ability to perform well with incomplete data (Stockman et al., 2023), typically a threshold below the completeness biases classical models such as ETAS (Seif et al., 2017).
-Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds (e.g.
-Mancini et al., 2022; Stockman et al., 2023), which we also facilitate in our datasets.
-Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law (Gutenberg & Richter, 1936), which states that the distribution of earthquake magnitudes follows an exponential probability density function
-$f_{GR}(m)=\beta e^{\beta(m-M_{c})}\ \ :m\geq M_{c}.$ (17)
+Where NPPs have demonstrated an ability to perform well with incomplete data [77], typically a threshold below the completeness biases classical models such as ETAS [67].
+Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds [e.g.
+40, 77], which we also facilitate in our datasets.
+Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law [17], which states that the distribution of earthquake magnitudes follows an exponential probability density function
+$f_{GR}(m)=\beta e^{-\beta(m-M_{c})}\ \ :m\geq M_{c}.$ (8)
 where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$.
-Histogram-based approaches, such as the simple Maximum Curvature method (Wiemer & Wyss, 2000) as well as many others (e.g.
-Herrmann & Marzocchi, 2021, and references therein), identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
-In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ (e.g.
-Schorlemmer & Woessner, 2008).
-During aftershock sequences, $M_{c}(t)$ can be very high (e.g., Agnew, 2015; Hainzl, 2016b) (See Figure 6a).
+Histogram-based approaches, such as the simple Maximum Curvature method [91] as well as many others [e.g.
+26, and references therein], identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
+In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ [e.g.
+65].
+During aftershock sequences, $M_{c}(t)$ can be very high [e.g., 1, 19] (See Figure 6a).
 Thresholding at the maximum value might remove too much data.
-Instead, modelers either omit particularly incomplete periods during training and testing (Kagan, 1991; Hainzl et al., 2008), model the incompleteness itself (Helmstetter et al., 2006; Werner et al., 2011; Omi et al., 2014; Hainzl, 2016a; b; Mizrahi et al., 2021; Hainzl, 2022), or accept known biases from disregarding this issue (Sornette & Werner, 2005).
-Spatially, catalogs are less complete farther from the seismic network (Mignan et al., 2011), so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
+Instead, modelers either omit particularly incomplete periods during training and testing [32, 21], model the incompleteness itself [25, 89, 55, 18, 19, 45, 20], or accept known biases from disregarding this issue [73].
+Spatially, catalogs are less complete farther from the seismic network [42], so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
 Figure 6: a) the June 10, 2016 Mw5.2 Borrego Springs earthquake and aftershocks, which occurred on the San Jacinto fault zone and is recorded in the White catalog.
@@
 c) An estimate of $M_{c}$ for gridded regions of the San Jacinto fault zone, using the raw White catalog.
-Appendix B Additional Benchmark Results
+Appendix B Additional Datasets
+Beyond the official EarthquakeNPP datasets, we include 3 further datasets that either provide additional scientific insight or continuity from previous benchmarking works.
+Synthetic ETAS Catalogs.
+We simulate a synthetic catalog using the ETAS model with parameters estimated from ComCat, at $M_{c}\ 2.5$, within the same California region.
+A second catalog emulates the time-varying data-missingness present in observational catalogs by removing events using the time-dependent formula from Page et al.
+[59],
+$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (9)
+where $M$ is the mainshock magnitude.
+Events below this threshold are removed using mainshocks of Mw 5.2 and above.
+The inclusion of these datasets allows us to test whether NPPs are inhibited by data missingness to the same extent that ETAS is.
+Deprecated Catalog of Japan.
+To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al.
+[5], however with a chronological train-test split and without removing any supposed outlier events.
+To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
+Table 3: Summary of additional datasets, including: magnitude threshold ($\mathbf{M}_{\mathbf{c}}$), number of training events, and number of testing events.
+The chronological partitioning of training, validation, and testing periods is also detailed.
+An auxiliary (burn-in) period begins from the "Start" date, followed by the respective starts of the training, validation, and testing periods.
+All dates are given as 00:$0$ UTC on January $1^{\text{st}}$, unless noted (* refers to 00:$0$ UTC on January $17^{\text{th}}$).
+Catalog $\mathbf{M}_{\mathbf{c}}$ Start-Train-Val-Test-End Train Events Test Events
+ETAS 2.5 1971-1981-1998-2007-2020 117,550 43,327
+ETAS_incomplete 2.5 1971-1981-1998-2007-2020 115,115 42,932
+Japan_Deprecated 2.5 1990-1992-2007-2011-2020 22,213 15,368
 Figures 7 and 8 report the temporal and spatial log-likelihood scores of all the benchmarked models on additional datasets.
@@
 Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
-Appendix C CSEP Consistency Tests
-Figure 9: Quantile-quantile plots showing the calibration of all daily ETAS forecasts on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
-By construction quantile scores over multiple periods should be uniformly distributed if the model is the data generator.
-Comparing quantile scores against standard uniform quantiles ($y=x$), highlights discrepancies between the observed data and the forecast.
-Pass rates of each test are indicated in the legend.
-The Kolmogorov-Smirnov statistic, quantifies the degree of difference to the uniform distribution.
+Appendix C Computational Efficiency
+C.1 Training
+Table 4 reports the training times for each model across all datasets.
+We ran all the NPP models using a HPC node with Nvidia Ampere GPU with 4x Nvidia A100 40GB SXM “Ampere” GPUs and AMD EPYC 7543P 32-Core Processor “Milan” CPU using torch==1.12.0 and cuda==11.3.
+Dataset # Training Events ETAS Deep-STPP AutoSTPP NSTPP SMASH DSTPP Poisson
+ComCat 79,037 08:59:04 00:15:35 01:34:09 3 days, 05:10:17 2:21:13 20:05:57 <1 second
+QTM_SaltonSea 44,042 07:28:28 00:26:46 01:45:34 2 days, 00:26:45 2:38:21 11:12:$0$ <1 second
+QTM_SanJac 18,664 00:32:40 00:09:31 00:37:03 1 day, 22:06:33 0:55:34 4:44:46 <1 second
+SCEDC_20 128,265 13:42:30 00:38:10 02:54:51 3 days, 02:20:40 4:22:16 1 day, 8:37:05 <1 second
+SCEDC_25 43,221 03:09:14 00:09:34 00:56:05 2 days, 16:33:55 1:24:07 10:59:28 <1 second
+SCEDC_30 12,426 00:42:25 00:02:44 00:16:01 1 day, 16:39:04 0:24:01 3:10:26 <1 second
+White 38,556 03:55:40 00:08:21 01:10:51 2 days, 01:03:57 1:46:17 9:48:47 <1 second
+Japan_Deprecated 22,213 06:09:08 00:13:45 01:02:07 2 days, 05:32:03 1:33:06 5:39:32 <1 second
+ETAS 117,550 00:33:25 00:15:24 01:10:22 3 days, 03:09:17 1:45:33 1 day, 1:27:44 <1 second
+ETAS_incomplete 115,115 00:35:14 00:15:29 01:09:43 3 days, 11:39:51 1:44:34 1 day, 2:28:42 <1 second
+[…]
+ETAS training scales $\mathcal{O}(n^{2})$ with the total number of events, since for every event a contribution to the intensity function is computed from a summation over all previous events.
+This scaling, coupled with the lack of parallelization in the current implementation, results in long training times for larger datasets.
+Poorer scaling will likely hinder ETAS if dataset sizes continue to grow in the future [76].
+Encouragingly, both DeepSTPP and AutoSTPP are significantly faster to train due to GPU acceleration and their use of a sliding window of the most recent $k=20$ events.
+While exact complexity analyses are not provided in Zhou et al.
+[97] or Zhou and Yu [96], we can infer that DeepSTPP likely scales as $\mathcal{O}(kn)$ since it benefits from a closed-form expression for the likelihood.
+AutoSTPP, though requiring automatic integration to compute the likelihood, still scales with $\mathcal{O}(kn)$ because the additional integration cost does not affect the overall scaling.
+NSTPP, on the other hand, incurs significant training costs, rendering it impractical for real-time forecasting.
+Unlike the sliding window mechanism used in DeepSTPP and AutoSTPP, NSTPP partitions the event sequence into fixed time intervals, leading to sequences that are much longer than the $k=20$ events used by the other models (as shown in Figure 11 of Chen et al.
+[5]).
+Furthermore, solving an ODE for each event time adds a significant computational burden, even with the use of their faster attentive CNF architecture.
+Whilst SMASH and DSTPP are built on the same backbone architecture, SMASH is much quicker to train than DSTPP—even faster than ETAS.
+This efficiency arises from its use of a single-step, normalization-free score-matching objective, which avoids the costly denoising and sampling loops required in diffusion-based training.
+SMASH directly learns the gradient of the log-density via pseudolikelihood estimation, enabling efficient GPU parallelization and bypassing the need for repeated evaluations over diffusion steps.
+In contrast, DSTPP simulates a sequential generative process over hundreds of intermediate steps per sample, significantly increasing computation and memory costs.
+C.2 Simulation
+Real-time earthquake forecasting and CSEP model evaluation require simulating many repeat sequences (at least 10,000 for adequate distributional coverage) over the forecasting horizon.
+While ETAS training scales as $\mathcal{O}(n^{2})$ with the number of training events, its simulation scales more efficiently at $\mathcal{O}(n\log n)$.
+This improved scaling is due to its equivalent formulation as a Hawkes branching process (see Section 2.2).
+Both Deep-STPP and AutoSTPP are also based on Hawkes processes, which theoretically allows for fast simulation.
+However, as these models currently only have an intensity function implementation, simulating events would require a slower thinning procedure [52], limiting their simulation efficiency.
+NSTPP too has slow simulation, since it must solve an ODE to sample a new event.
+Due to the slow simulation of DeepSTPP, AutoSTPP and NSTPP, they are not evaluated using the CSEP generative evaluation metrics.
+Appendix D Analysis of Likelihood Scores
+(a) ComCat
+(b) SCEDC
+(c) SanJac
+(d) SaltonSea
+(e) WHITE
+Figure 9: Cumulative information gain (IG) plots for the temporal performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
+(a) ComCat
+(b) SCEDC
+(c) SanJac
+(d) SaltonSea
+(e) WHITE
+Figure 10: Cumulative information gain (IG) plots for the spatial performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
+Appendix E CSEP Consistency Tests
+E.1 Number (Temporal) Test
+The number test evaluates the temporal component of the forecast by checking the consistency of the forecasted number of events, $N$ with those observed in the forecast horizon, $N_{\text{obs}}$.
+Upper and lower quantiles are estimated using the empirical cumulative distribution from the repeat simulations, $F_{N}$,
+$\delta_{1}$ $=\mathbb{P}(N\geq N_{\text{obs}})=1-F_{N}(N_{\text{obs}}-1)$ (10)
+$\delta_{2}$ $=\mathbb{P}(N\leq N_{\text{obs}})=F_{N}(N_{\text{obs}}).$ (11)
+E.2 Pseudo-Likelihood Test
+The pseudo-likelihood test evaluates the compatibility of a forecast with an observed catalog using an approximation to the space-time point process likelihood.
+The test statistic is based on the pseudo-log-likelihood:
+$\hat{L}_{\text{obs}}=\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}_{s}(k_{i})-\bar{N},$ (12)
+where $\hat{\lambda}_{s}(k_{i})$ is the approximate rate density in the spatial cell of the $i^{\text{th}}$ event, and $\bar{N}$ is the expected number of events.
+Each forecast simulation $j$ provides a test statistic
+$\hat{L}_{j}=\sum_{i=1}^{N_{j}}\log\hat{\lambda}_{s}(k_{ij})-\bar{N},$ (13)
+which is used to build the empirical cumulative distribution $F_{L}$. The quantile score is then computed as
+$\gamma_{L}=\mathbb{P}(\hat{L}_{j}\leq\hat{L}_{\text{obs}})=F_{L}(\hat{L}_{\text{obs}}).$ (14)
+E.3 Spatial Test
+To evaluate the spatial component of the forecast, a test statistic aggregates the forecasted rates of earthquakes over a regular grid,
+$S=\left[\sum_{i=1}^{N}\log\hat{\lambda}(k_{i})\right]N^{-1},$ (15)
+where $\hat{\lambda}(k_{i})$ is the approximate rate in the cell $k$ where the $i^{th}$ event is located. Upper and lower quantiles are estimated by comparing the observed statistic
+$S_{\text{obs}}=\left[\sum_{i=1}^{N_{\text{obs}}}\log\hat{\lambda}(k_{i})\right]N_{\text{obs}}^{-1},$ (16)
+with the empirical cumulative distribution of $S$ using the repeat simulations, $F_{S}$
+$\gamma_{s}=\mathbb{P}(S\leq S_{\text{obs}})=F_{S}(S_{\text{obs}}).$ (17)
+The grid is constructed from $\{0.1\degree,0.05\degree,0.01\degree\}$ squares for ComCat, SCEDC and {QTM_Salton_Sea, QTM_SanJac, White} respectively.
+E.4 Magnitude Test
+To evaluate the earthquake magnitude component of the forecast, a test statistic compares the histogram of a forecast’s magnitudes $\Lambda^{(m)}$, against the mean histogram over all forecasts $\bar{\Lambda}^{(m)}$,
+$D=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda^{(m)}(k)+1\right]\right)^{2},$ (18)
+where $\Lambda^{(m)}(k)$ and $\bar{\Lambda}^{(m)}(k)$ are the counts in the $k^{th}$ bin of the forecast and mean histograms, normalised to have the same total counts as the observed catalog.
+Upper and lower quantiles are estimated by comparing the observed statistic
+$D_{\text{obs}}=\sum_{k}\left(\log\left[\bar{\Lambda}^{(m)}(k)+1\right]-\log\left[\Lambda_{\text{obs}}^{(m)}(k)+1\right]\right)^{2},$ (19)
+with the empirical distribution of $D$ using the repeat simulations, $F_{D}$
+$\gamma_{m}=\mathbb{P}(D\leq D_{\text{obs}})=F_{D}(D_{\text{obs}}).$ (20)
+Histogram bins of size $\delta_{m}=0.1$ are used across all datasets.
+(a) Number test
+(b) Spatial test
+(c) Magnitude test
+Figure 11: CSEP consistency tests on the ETAS model for the first day ($01/01/2014$) of the testing period in the SCEDC catalog.
+A total of 10,000 simulations are generated to compute empirical distributions of the test statistics for each of the three consistency tests: (a) Number test, (b) Spatial test, and (c) Magnitude test.
+The test fails if the observed statistic falls within the rejection region (red), defined by the $0.05$ and $0.95$ quantiles of the distribution.
+E.5 Evaluating Multiple Forecasting Periods
+Savran et al.
+[63] describe how to assess a model’s performance across the multiple days in the testing period (Figure 12).
+By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator [15].
+Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast.
+Additional statements can be made about over-prediction or under-prediction of each test statistic (quantile curves above/bellow y=x respectively).
+The Kolmogorov-Smirnov (KS) statistic then quantifies the degree of difference to the uniform distribution for each of the tests.
+Figure 12: Daily number forecasts from SMASH on the ComCat dataset.
+(a) Forecasted daily distributions for the number of earthquakes, with green lines indicating days where the observed count falls within the 95% forecast interval, and red lines where the forecast fails.
+Observed values are marked with dot sizes proportional to the number of earthquakes.
+(b) Quantile scores from the number test for each day, with red markers indicating failed forecasts.
+Marker size reflects the number of earthquakes observed on that day.
+(c) Temporal evolution of observed earthquakes during the testing period, with event magnitudes represented by marker size.
+(d) Histogram of quantile scores from the number test.
+Under ideal calibration, scores should follow a uniform distribution.
+Red bars indicate failed forecasts, and the Kolmogorov–Smirnov (KS) statistic quantifies deviation from uniformity.
+Appendix F Further Dataset Figures
+F.1 ComCat
+Figure 13: Times and magnitudes of events in the ComCat dataset (with key events labeled).
+The size of the points are plotted on a log scale corresponding to Mw.
+Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
+Figure 14: Locations of events in the ComCat dataset, labeled by their partition into auxiliary, training, validation and testing periods.
+F.2 SCEDC
+Figure 15: Times and magnitudes of events in the SCEDC dataset (with key events labeled).
+The size of the points are plotted on a log scale corresponding to Mw.
+Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
+Figure 16: Locations of events in the SCEDC dataset, labeled by their partition into auxiliary, training, validation and testing periods.
+F.3 White
+Figure 17: Times and magnitudes of events in the White dataset (with key events labeled).
+The size of the points are plotted on a log scale corresponding to Mw.
+Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
+Figure 18: Locations of events in the White dataset, labeled by their partition into auxiliary, training, validation and testing periods.
+F.4 QTM_SanJac
+Figure 19: Times and magnitudes of events in the QTM_SanJac dataset.
+The size of the points are plotted on a log scale corresponding to Mw.
+Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
+F.5 QTM_SaltonSea
+Figure 20: Times and magnitudes of events in the QTM_SaltonSea dataset.
+The size of the points are plotted on a log scale corresponding to Mw.
+Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
+Figure 21: Locations of events in the QTM_SanJac and QTM_SaltonSea datasets, labeled by their partition into auxiliary, training, validation and testing periods.
+Appendix G Error Distributions & Next-event metrics
+Figure 22: The distribution of errors ($Y_{\text{obs}}-Y_{\text{pred}}$) for the Normal(0,1)(0,1), Exponential(1)(1), and Pareto(2)(2) distributions.
+Maximum likelihood estimation is used to fit Normal and Laplace distributions to each error histogram.
+Normal errors (Normal $\times$ Normal) are best approximated by the Root Mean Square Error (RMSE), while Laplacian errors (Exponential $\times$ Exponential) are best approximated by the Mean Absolute Error (MAE).
+However, neither RMSE nor MAE effectively capture the errors for the heavy-tailed Pareto distribution.

### v3

@@
 Additionally, initial earthquake forecasting benchmarks fail to compare NPPs with state-of-the-art forecasting models commonly used in seismology.
-To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by an implementation of the state-of-the-art forecasting model: ETAS.
+To address these gaps, we introduce EarthquakeNPP: a benchmarking platform that curates and standardizes existing public resources: globally available earthquake catalogs, the ETAS model, and evaluation protocols from the seismology community.
 The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation.
@@
 Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes.
-While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model [53, 54], despite significant growth in available data [79, 72, 62, 90, 51, 80, 50].
-In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility [9, 56, 69, 31, 5, 97, 96].
+While model development is ongoing in the seismology community, recent progress has relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model (Ogata, 1988; 1998).
+This continued reliance on a low-dimensional parametric framework stands in contrast to the substantial growth in available earthquake data (Takanami et al., 2003; Shelly, 2017; Ross et al., 2019; White et al., 2019; Mousavi et al., 2020; Tan et al., 2021; Mousavi and Beroza, 2023).
+In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019; Jia and Benson, 2019; Chen et al., 2021; Zhou et al., 2022; Zhou and Yu, 2024).
 While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community.
-The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
+The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and does not compare against state-of-the-art models like ETAS.
 Here, we introduce EarthquakeNPP: a curated collection of datasets designed for benchmarking NPP models in earthquake forecasting, accompanied by a state-of-the-art benchmark model.
@@
 In addition to the standard log-likelihood metric common in the NPP literature, the platform incorporates the generative evaluation procedures used in seismology for more rigorous benchmarking.
-This ensures that future NPPs (and other models such as time series approaches [88] and Bayesian point processes [68]) can have direct relevance to seismological stakeholders.
-All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP/tree/main.
+This ensures that future NPPs (and other models such as time series approaches (Wang et al., 2017) and Bayesian point processes (Serafini et al., 2023)) can have direct relevance to seismological stakeholders.
+All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP.
 1.1 Related Work
-Benchmarking by the NPP Community.
+1.1.1 Benchmarking by the NPP Community
 Chen et al.
-[5] introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S.
+(2021) introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S.
 Geological Survey, focusing on Japan from 1990 to 2020.
 They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset.
-They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake [49], the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0.
+They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers".
+However, this period includes the 2011 Tohoku earthquake (Mori et al., 2011), the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0.
 This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact.
-The dataset segments are divided for training, testing, and validation.
-Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern.
-This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering [13] - a form of data leakage.
-Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards in time.
+The dataset is partitioned into training, testing, and validation segments.
+Rather than following a chronological split that would reflect operational forecasting, the segments are assigned in an alternating pattern.
+This design introduces data leakage, as it misrepresents a realistic forecasting setup and artificially inflates performance measures due to the nature of earthquake triggering (Freed, 2005).
+Specifically, because the model is evaluated on windows that immediately precede its training windows, it can exploit backward-in-time causal dependencies.
+Section B.2 quantifies the resulting performance inflation, expressed in terms of information gain.
 Although earthquakes with magnitudes above 2.5 are considered by Chen et al.
-[5], following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset.
+(2021), following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset.
 For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets.
@@
 Chen et al.
-[5] benchmark their model against another spatio-temporal model, Neural Jump SDEs [31], and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark.
-Subsequent papers adopting this benchmark [97, 94, 96] similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
-Benchmarking by the Seismology Community.
-Model comparison has been crucial in the development of earthquake forecasting models since their inception [34, 53].
-The Collaboratory for the Study of Earthquake Predictability (CSEP) [41, 66, 64, 30] [https://cseptesting.org/ ] aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally.
+(2021) benchmark their model against another spatio-temporal model, Neural Jump SDEs (Jia and Benson, 2019), and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark.
+Subsequent papers adopting this benchmark (Zhou et al., 2022; Yuan et al., 2023; Zhou and Yu, 2024) similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
+1.1.2 Benchmarking by the Seismology Community.
+Model comparison has been crucial in the development of earthquake forecasting models since their inception (Kagan and Knopoff, 1987; Ogata, 1988).
+The Collaboratory for the Study of Earthquake Predictability (CSEP) (Michael and Werner, 2018; Schorlemmer et al., 2018; Savran et al., 2022; Iturrieta et al., 2024) (https://cseptesting.org/ ) aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally.
 CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day).
 These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times.
-The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
+The simulation-based approach allows the inclusion of generative models that do not output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
 Two existing works benchmark NPPs for earthquake forecasting within the seismology community.
 The first by Dascher-Cousineau et al.
-[8] extends a temporal-only NPP from Shchur et al.
-[69] to include earthquake magnitudes.
+(2023) extends a temporal-only NPP from Shchur et al.
+(2019) to include earthquake magnitudes.
 The second by Stockman et al.
-[77] extends another temporal-only model by Omi et al.
-[56] to target larger magnitude events.
+(2023) extends another temporal-only model by Omi et al.
+(2019a) to target larger magnitude events.
 Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline.
@@
 2.1 Spatio-Temporal Point Processes
-A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$.
+A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\subset\mathbb{R}^{+}$.
 This process is typically defined by a non-negative conditional intensity function
@@
 $p(t,\mathbf{x}|\mathcal{H}_{t_{i}})=\lambda(t,\mathbf{x}|\mathcal{H}_{t_{i}})\cdot\exp\left(-\int_{t_{i}}^{t}\int_{\mathcal{S}}\lambda(s,\mathbf{z}|\mathcal{H}_{s})d\mathbf{z}ds\right).$ (2)
-Most models specify the conditional intensity function, though some [e.g.
-69, 5, 94] directly model this pdf.
+Most models specify the conditional intensity function, though some (e.g.
+Shchur et al., 2019; Chen et al., 2021; Yuan et al., 2023) directly model this pdf.
 Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
@@
 For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric.
-This is consistent with a wealth of literature in the seismology community [see 95, and references therein] as well as the wider general point process literature [7], which now includes neural point processes [70].
+This is consistent with a wealth of literature in the seismology community (see Zechar et al., 2010, and references therein) as well as the wider general point process literature (Daley and Vere-Jones, 2004), which now includes neural point processes (Shchur et al., 2021).
 The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident.
 Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function.
-Comparing the difference in mean log-likelihood per event provides the information gain from one model to another [7].
-Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting [47] and established benchmarking experiments (CSEP) [81, 60].
-The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time [33].
-Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin — a delay that could miss critical, potentially damaging events.
+Comparing the difference in mean log-likelihood per event provides the information gain from one model to another (Daley and Vere-Jones, 2004).
+Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting (Mizrahi et al., 2024a) and established benchmarking experiments (CSEP) (Taroni et al., 2018; Rhoades et al., 2018).
+The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time (Kagan, 1994).
+Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin - a delay that could miss critical, potentially damaging events.
 2.2 ETAS
-The Epidemic Type Aftershock Sequence (ETAS) model [54] is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space.
-It has been adopted for operational earthquake forecasting by government agencies in California [44], New-Zealand [6], Italy [74], Japan [57] and Switzerland [48], and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments [e.g.
-93, 60, 81, 4, 38, 39, 40].
+The Epidemic Type Aftershock Sequence (ETAS) model (Ogata, 1998) is a spatio-temporal Hawkes process Hawkes (1971); Siviero et al.
+(2024); Bernabeu and Mateu (2025) which models how earthquakes cluster in time and space.
+It has been adopted for operational earthquake forecasting by government agencies in California (Milner et al., 2020), New-Zealand (Christophersen et al., 2017), Italy (Spassiani et al., 2023), Japan (Omi et al., 2019b) and Switzerland (Mizrahi et al., 2024b), and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments (e.g.
+Woessner et al., 2011; Rhoades et al., 2018; Taroni et al., 2018; Cattania et al., 2018; Mancini et al., 2019; 2020; 2022).
 The general formulation of the model is
@@
 This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g.
-via dynamic wave triggering [3] and static stress triggering [16, 39].
+via dynamic wave triggering (Brodsky and van der Elst, 2014) and static stress triggering (Gomberg, 2018; Mancini et al., 2020).
 Although ETAS can be fit by maximizing the log-likelihood function directly, parameter estimation is typically performed by simultaneously estimating the branching structure $\mathbf{B}$.
-Veen and Schoenberg [87] developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
+Veen and Schoenberg (2008) developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
 $\mathbf{\theta}^{(k+1)}=\operatorname*{arg\,max}_{\mathbf{\theta}}\mathbb{E}_{\mathbf{B}\sim p(\cdot|\mathcal{H}_{T_{1}},\mathbf{\theta}^{(k)})}\left[\log p(\mathcal{H}_{T_{1}},\mathbf{B}|\mathbf{\theta})\right].$ (5)
 This avoids the need to numerically approximate the integral term in the likelihood, provides more stability during estimation, and simultaneously distinguishes background events from triggered events.
+The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. (2022). It defines the triggering kernel as
+$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{c})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{c})}\right)^{1+\rho}},$ (6)
+where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$.
+This triggering kernel is derived from statistical distributions found through decades of observational studies (Utsu and Seki, 1955; Utsu, 1970; Utsu et al., 1995) and several of the learnable parameters have been linked to physical properties of the earthquake rupture process (Utsu et al., 1995; Ide, 2013).
+Despite its widespread use, it is commonly accepted that ETAS is a misspecified model of seismicity.
+By construction, ETAS describes only self-exciting triggering behaviour and therefore cannot capture inhibitory effects or stress relaxation processes, such as those represented by stress-release models (Zheng and Vere-Jones, 1991; Xiaogu and Vere-Jones, 1994; Bebbington and Harte, 2003), or by models based on elastostatic stress transfer and Coulomb rate-and-state friction (Dieterich, 1994).
+In addition, foreshock activity has been observed to deviate from ETAS assumptions, both spatially and temporally (McGuire et al., 2005; Brodsky, 2011; Lippiello et al., 2012; Ogata and Katsura, 2014).
+Finally, to simplify inference, ETAS typically assumes isotropic spatial triggering kernels, despite observational evidence for anisotropic and fault-aligned aftershock distributions (Page and van der Elst, 2022).
+Together, these limitations motivate the exploration of more flexible modelling frameworks capable of capturing richer spatio-temporal structure in earthquake sequences.
+2.3 Neural Point Processes
+Neural point processes (NPPs) have emerged in recent years within the machine learning literature as flexible alternatives to classical parametric point process models.
+Their central motivation is to replace restrictive, hand-crafted parametric forms with neural network based components that can learn complex, non-linear dependencies directly from data.
+This makes them particularly appealing for earthquake forecasting to overcome the known limitations of the ETAS model discussed in Section 2.2.
+Early developments focused on temporal point processes (Shchur et al., 2021).
+Du et al.
+(2016) introduced the use of recurrent neural networks (RNNs) to encode the event history into a fixed-dimensional latent state, replacing explicit summation over past events with a learned representation of temporal dependence.
+Subsequent work explored alternative sequence encoders, including LSTMs (Mei and Eisner, 2017) and Transformers (Zuo et al., 2020), alongside a variety of decoding strategies for modelling the conditional intensity or inter-event time distribution (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019).
+In most cases, model parameters are learned by maximising the log-likelihood of observed event sequences, although alternative training objectives have also been proposed (Xiao et al., 2017; Li et al., 2018).
+These temporal formulations were later extended to spatio-temporal settings (Mukherjee et al., 2025) by incorporating event locations into the history encoder and introducing flexible decoders for the conditional spatial distribution of future events.
+Existing spatio-temporal NPPs can be broadly grouped into three modelling classes.
+The first class extends Hawkes-type formulations by replacing parametric triggering kernels with neural network based influence functions, allowing non-stationary and history-dependent excitation (Zhou et al., 2022; Dong et al., 2022; Zhou and Yu, 2024).
+A second class models event dynamics in continuous time using neural ordinary differential equations, jointly evolving latent temporal states and spatial distributions (Jia and Benson, 2019; Chen et al., 2020).
+A third, more recent class adopts fully generative approaches based on diffusion or score matching, learning the joint spatio-temporal distribution of events without explicitly parameterising an intensity function (Yuan et al., 2023; Li et al., 2023; Lüdke et al., 2024).
+These approaches differ substantially in their computational cost, interpretability, and suitability for simulation and likelihood-based evaluation; see Mukherjee et al.
+(2025) for a detailed discussion.
 Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP.
 Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans.
-In red is a fault map from the GEM Global Active Faults Database [78].
-The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. [46]. It defines the triggering kernel as
-$g(t,r^{2},m)=\frac{e^{{-t}/{\tau}}\cdot k\cdot e^{a(m-M_{c})}}{(t+c)^{1+\omega}\cdot\left(r^{2}+d\cdot e^{\gamma(m-M_{c})}\right)^{1+\rho}},$ (6)
-where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$.
-This triggering kernel is derived from statistical distributions found through decades of observational studies [84, 83, 85] and several of the learnable parameters have been linked to physical properties of the earthquake rupture process [85, 29].
+In red is a fault map from the GEM Global Active Faults Database (Styron and Pagani, 2020).
 3 EarthquakeNPP Datasets
 The EarthquakeNPP datasets (Figure 1) encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021.
-California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms [14, 11, 12].
-It encompasses the San Andreas fault plate boundary system [99] and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
-Notebooks to access and preprocess these public datasets along with the associated benchmarking experiment are publicly accessible at https://github.com/ss15859/EarthquakeNPP/tree/main, accompanied by more detailed documentation for each dataset.
-A summary of how earthquake datasets are generated, along with the associated challenges of using earthquake catalog data can be found in Appendix A.
-Table 1 provides a short summary of each EarthquakeNPP dataset.
+California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms (Gerstenberger et al., 2004; Field, 2007; Field et al., 2021).
+It encompasses the San Andreas fault plate boundary system (Zoback et al., 1987) and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
 Table 1: Summary of EarthquakeNPP datasets, including: region, dataset development, magnitude threshold ($\mathbf{M}_{\mathbf{c}}$), number of training (combined with validation) events, and number of testing events.
@@
 Its contributing seismic networks have produced the Advanced National Seismic System (ANSS) Comprehensive Catalog of Earthquake Events and Products.
-The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California [28] since the Caltech Seismological Laboratory began routine operations in 1932.
+The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California (Hutton et al., 2010) since the Caltech Seismological Laboratory began routine operations in 1932.
 Significant network improvements since the 1970s and 1980s reduced the catalog completeness from Mw 3.25 to Mw 1.8.
 White et al.
-[90] created an enhanced catalog focusing on the San Jacinto fault region, using a dense seismic network in Southern California.
-This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion [90].
+(2019) created an enhanced catalog focusing on the San Jacinto fault region, using a dense seismic network in Southern California.
+This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion (White et al., 2019).
 Using data collected by the SCSN, Ross et al.
-[62] generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
+(2019) generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
 $\mathbf{M}_{\mathbf{c}}$ Mw 2.5 SCEDC_20: Mw 2.0,
@@
 Purpose Example of data currently in use for operational
-forecasting (USGS utilizes ComCat in aftershock forecasts they release to the public.)
-Three magnitude thresholds (Mw 2.0, 2.5, 3.0) explore the effect of truncation on forecasting model performance.
+forecasting (USGS utilizes ComCat in aftershock forecasts they release to the public).
+Three magnitude thresholds (Mw 2.0, 2.5, 3.0 for SCEDC_20, SCEDC_25, SCEDC_30) explore the effect of truncation on forecasting model performance.
 To explore if newly detected low magnitude earthquakes contain additional predictive information.
 To explore if newly detected low magnitude earthquakes contain additional predictive information (with different detection methodology).
+A central challenge when working with earthquake catalogs is data missingness, referred to in seismology as catalog incompleteness (Mignan and Woessner, 2012).
+Earthquakes are assumed to be only fully detected above a time- and region-dependent completeness magnitude $M_{c}$, which reflects limitations of the seismic network and changes in detection capability over time.
+Ignoring this incompleteness can introduce substantial bias into both model fitting and evaluation (Sornette and Werner, 2005), particularly for methods that rely on smaller magnitude events.
+All EarthquakeNPP datasets are constructed from publicly available raw catalogs provided by their respective data centres.
+To enable a consistent and realistic retrospective forecasting experiment, the raw data is preprocessed by restricting each dataset to a target spatial region, truncating events above a dataset-specific magnitude threshold $M_{\text{cut}}\geq M_{c}$ (e.g., Mignan et al., 2011; Mignan and Woessner, 2012) and removing duplicate locations.
+Notebooks to access and preprocess the datasets, along with the associated benchmarking experiments, are publicly available at https://github.com/ss15859/EarthquakeNPP, accompanied by detailed dataset documentation.
+A more in-depth discussion of earthquake catalog generation, completeness, and preprocessing choices is provided in Appendix A.
+Table 1 summarises the key characteristics of each EarthquakeNPP dataset.
 4 Benchmarking Experiment
-We use EarthquakeNPP to benchmark five spatio-temporal Neural Point Processes (NPPs), each with prior positive claims in earthquake forecasting, against the ETAS model described in Section 2.2.
-Neural Spatio-Temporal Point Process (NSTPP) [5]: A probability density function (pdf)-based NPP that parametrizes the spatial pdf with continuous-time normalizing flows (CNFs).
-We evaluate their Attentive CNF model due to its superior computational efficiency and overall performance compared to the Jump CNF model [5].
-Deep Spatio-Temporal Point Process (DeepSTPP) [97]: A conditional intensity function-based NPP that constructs a non-parametric space-time intensity function driven by a deep latent process.
+We use EarthquakeNPP to benchmark five spatio-temporal Neural Point Processes (NPPs) against the ETAS model described in Section 2.2.
+Each of these NPPs has prior positive claims in earthquake forecasting, yet lacks performance comparison with the ETAS model.
+Neural Spatio-Temporal Point Process (NSTPP) (Chen et al., 2021): A probability density function (pdf)-based NPP that parametrizes the spatial pdf with continuous-time normalizing flows (CNFs).
+We evaluate their Attentive CNF model due to its superior computational efficiency and overall performance compared to the Jump CNF model (Chen et al., 2021).
+Deep Spatio-Temporal Point Process (DeepSTPP) (Zhou et al., 2022): A conditional intensity function-based NPP that constructs a non-parametric space-time intensity function driven by a deep latent process.
 This model features a closed-form intensity function, eliminating the need for numerical approximations.
-Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) [96]: A conditional intensity function-based NPP that jointly models the 3D space-time integral of the intensity and its derivative (the intensity function) using a dual network approach.
-Spatio-temporal Diffusion Point Process (DSTPP) [94]: A generative NPP that does not have a likelihood function. DSTPP employs diffusion models to capture complex spatio-temporal dynamics.
-Score Matching-based Pseudolikelihood Estimation of Neural Marked Spatio-Temporal Point Process (SMASH)[36]: A generative NPP that also lacks a likelihood function.
+Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) (Zhou and Yu, 2024): A conditional intensity function-based NPP that jointly models the 3D space-time integral of the intensity and its derivative (the intensity function) using a dual network approach.
+Spatio-temporal Diffusion Point Process (DSTPP) (Yuan et al., 2023): A generative NPP that does not have a likelihood function.
+DSTPP employs diffusion models to capture complex spatio-temporal dynamics.
+Score Matching-based Pseudolikelihood Estimation of Neural Marked Spatio-Temporal Point Process (SMASH)(Li et al., 2023): A generative NPP that also lacks a likelihood function.
 SMASH adopts a normalization-free objective by estimating the pseudolikelihood of marked STPPs through score-matching.
+Appendix D provides details on the computational cost of training and inference for all the models tested.
 4.1 Likelihood Evaluation
-Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development [53, 23].
-Other traditional next-event metrics like Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are misleading for earthquake forecasting [27], as earthquake occurrence follows power law distributions [33, 10] that are heavy-tailed, making the errors non-Gaussian and non-Laplacian — contrary to the assumptions underlying RMSE and MAE (see Section G).
+Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development (Ogata, 1988; Harte, 2015).
+Other traditional next-event metrics like Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are misleading for earthquake forecasting (Hodson, 2022), as earthquake occurrence follows power law distributions (Kagan, 1994; Felzer and Brodsky, 2006) that are heavy-tailed, making the errors non-Gaussian and non-Laplacian, contrary to the assumptions underlying RMSE and MAE (see Section I).
+Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
+SCEDC_20, SCEDC_25 and SCEDC_30 correspond to magnitude thresholds (Mw 2.0, 2.5, 3.0) of the SCEDC dataset.
+Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
+Figure 3: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
+SCEDC_20, SCEDC_25 and SCEDC_30 correspond to magnitude thresholds (Mw 2.0, 2.5, 3.0) of the SCEDC dataset.
+Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
 For the three models with valid likelihood functions (NSTPP, DeepSTPP, and AutoSTPP), we present the mean log-likelihood scores in Figures 2 and 3.
@@
 Nevertheless, its likelihood is evaluated on the same data as the other models.
-The definition of the ETAS model (equation 4) specifies how the magnitudes of earthquakes in the history contribute towards the intensity function.
-This earthquake magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modeling choices beyond the scope of this work.
+To ensure that fitting ETAS on both the training and validation windows does not bias the comparison, we also tested an alternative configuration where ETAS was trained only on the training window.
+As shown in Appendix C, ETAS performance remains effectively unchanged under this setup.
+The ETAS formulation (Equation 4) also specifies how the magnitudes of prior earthquakes contribute to the conditional intensity; this magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modelling choices beyond the scope of this work.
 The ETAS model consistently achieves the highest temporal log-likelihood, with NPPs performing comparably or, in some cases, marginally better, except at the higher magnitude thresholds of the SCEDC catalog.
 Among the NPPs, AutoSTPP and NSTPP perform well across several datasets, though their performance is more variable than that of DeepSTPP, which demonstrates consistent, comparable performance to ETAS.
-Closer examination of model performance over time (Figure 9) reveals that relative performance to ETAS is poorest during large earthquake sequences.
+Differences in Poisson performance across Figures 2 and 3 are driven by variations in clustering strength, with weakly clustered catalogs appearing nearly Poisson-like and strongly clustered catalogs exhibiting larger departures.
+Closer analysis of model performance over time (see Section E) reveals that relative performance to ETAS is poorest during large earthquake sequences.
 This is likely due to ETAS leveraging the magnitude feature of the data, which enables it to handle these sequences effectively.
@@
 During these periods, ETAS models the background with a constant rate, while the NPPs improve upon this by capturing the non-stationary nature of the background data.
-This effect is most pronounced in the ComCat dataset, which includes more complex physical processes, such as those near the Mendocino Triple Junction [24].
 The improved relative temporal performance of all NPPs compared to ETAS, particularly when the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, indicates that low magnitude earthquakes provide valuable predictive information for NPPs.
 ETAS consistently outperforms all NPPs in spatial log-likelihood.
-As with temporal performance, relative performance to ETAS is weakest during large earthquake periods, likely due to the absence of a magnitude feature in the NPPs.
-AutoSTPP achieves the highest spatial log-likelihood, attributed to its ability to capture anisotropic Hawkes kernels (see Figure 2 of Zhou and Yu [96]), which are commonly observed in earthquake data [58].
-Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
-Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
-Figure 3: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
-Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
+Further analysis of model performance over space (see Section E) shows relative performance to ETAS is weakest in the most active and clustered areas (see Figures 12 and 13), likely due to the absence of a magnitude feature in the NPPs.
+However, NPPs tend to perform more competitively in regions characterised by spatially complex or diffuse seismicity.
+AutoSTPP achieves the highest spatial log-likelihood, attributed to its ability to capture anisotropic Hawkes kernels (see Figure 2 of Zhou and Yu (2024)), which are commonly observed in earthquake data (Page and van der Elst, 2022).
 4.2 CSEP Consistency Tests
@@
 In this procedure, a model generates 24-hour forecasts through 10,000 repeated simulations of earthquake sequences at the beginning of each day in the testing period.
-This approach mirrors how earthquake forecasts are produced in operational settings [86].
+This approach mirrors how earthquake forecasts are produced in operational settings (van der Elst et al., 2022).
 Models are then evaluated by comparing the observed sequence with the distribution of forecasts generated by the simulations.
@@
 We apply this evaluation procedure to the two generative NPPs (DSTPP and SMASH) alongside ETAS (Table 2) and present a case study on the 2010 M7.2 El Mayor-Cucapah earthquake, using the forecasts from these models (Figure 4).
-Appendix E provides an introduction to the CSEP consistency tests, with further details found at https://cseptesting.org/.
-Table 2: CSEP consistency tests evaluate the calibration of daily forecasts from three models (ETAS, SMASH, DSTPP) on EarthquakeNPP datasets.
-A test is performed at the $\alpha=0.05$ significance level on each day in the testing period.
-The pass rate indicates the proportion of testing days with non-rejected hypotheses.
-If the model is the data generator, quantile scores should be uniformly distributed.
-The KS-Statistic quantifies deviation from uniformity (see Appendix E).
-ETAS is the only model that forecasts earthquake magnitudes, so is the only model evaluated with the magnitude test.
-Dataset Model Number Test Spatial Test Pseudo Likelihood Test Magnitude Test
-Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat.
-ComCat ETAS 85.1% 0.222 92.0% 0.029 97.6% 0.128 93.8% 0.113
-SMASH 72.4% 0.212 68.6% 0.217 87.6% 0.134
-DSTPP 70.4% 0.116 88.6% 0.070 86.3% 0.138
-SCEDC ETAS 81.7% 0.347 88.3% 0.119 95.9% 0.233 90.0% 0.256
-SMASH 59.9% 0.602 51.1% 0.471 68.0% 0.611
-DSTPP 0.0% 0.840 6.1% 0.935 0.8% 0.988
-QTM_SanJac ETAS 88.8% 0.151 91.7% 0.095 96.6% 0.123 94.8% 0.076
-SMASH 67.8% 0.290 55.6% 0.385 53.4% 0.584
-DSTPP 78.4% 0.110 85.7% 0.240 85.3% 0.136
-QTM_SaltonSea ETAS 77.8% 0.210 90.9% 0.206 96.4% 0.119 90.6% 0.136
-SMASH 58.6% 0.486 53.6% 0.371 73.7% 0.451
-DSTPP 69.6% 0.154 88.8% 0.136 85.6% 0.127
-White ETAS 83.4% 0.167 86.6% 0.225 90.8% 0.233 88.8% 0.052
-SMASH 61.3% 0.274 84.5% 0.150 67.7% 0.246
-DSTPP 0.0% 0.987 30.9% 0.691 32.3% 0.892
-ETAS consistently performs best across all datasets and tests.
-It achieves the highest pass rates and lowest KS statistics, indicating strong calibration and reliability.
-SMASH shows moderate performance, often outperforming DSTPP but trailing ETAS.
-Its results vary more across datasets and tests, with occasional strengths (e.g.
-in White for spatial KS).
-DSTPP generally performs worse, with lower pass rates and higher KS statistics, especially for the SCEDC and White datasets.
-However, it achieves relatively good spatial calibration in some cases (e.g., ComCat).
+Appendix F provides an introduction to the CSEP consistency tests, with further details found at https://cseptesting.org/, and Appendix G provides further analysis on the simulated forecasts.
 Figure 4: Forecasts from ETAS, SMASH, and DSTPP during the 2010 M7.2 El Mayor-Cucapah earthquake contained in the ComCat dataset.
@@
 ETAS and SMASH begin to recover after the first week, whereas DSTPP continues to systematically underpredict event counts throughout the sequence.
+Table 2: CSEP consistency tests evaluate the calibration of daily forecasts from three models (ETAS, SMASH, DSTPP) on EarthquakeNPP datasets.
+A test is performed at the $\alpha=0.05$ significance level on each day in the testing period.
+The pass rate indicates the proportion of testing days with non-rejected hypotheses.
+If the model is the data generator, quantile scores should be uniformly distributed.
+The KS-Statistic quantifies deviation from uniformity (see Appendix F).
+ETAS is the only model that forecasts earthquake magnitudes, so is the only model evaluated with the magnitude test.
+Dataset Model Number Test Spatial Test Pseudo Likelihood Test Magnitude Test
+Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat. Pass Rate KS-Stat.
+ComCat ETAS 95.8% 0.222 92.0% 0.029 97.6% 0.128 93.8% 0.113
+SMASH 86.2% 0.212 68.6% 0.217 87.6% 0.134
+DSTPP 86.7% 0.116 88.6% 0.070 86.3% 0.138
+SCEDC ETAS 92.6% 0.347 88.3% 0.119 95.9% 0.233 90.0% 0.256
+SMASH 69.6% 0.602 51.1% 0.471 68.0% 0.611
+DSTPP 27.4% 0.840 6.1% 0.935 0.8% 0.988
+QTM_SanJac ETAS 95.4% 0.151 91.7% 0.095 96.6% 0.123 94.8% 0.076
+SMASH 81.7% 0.290 55.6% 0.385 53.4% 0.584
+DSTPP 85.7% 0.110 85.7% 0.240 85.3% 0.136
+QTM_SaltonSea ETAS 93.6% 0.210 90.9% 0.206 96.4% 0.119 90.6% 0.136
+SMASH 75.8% 0.486 53.6% 0.371 73.7% 0.451
+DSTPP 87.7% 0.154 88.8% 0.136 85.6% 0.127
+White ETAS 88.3% 0.167 86.6% 0.225 90.8% 0.233 88.8% 0.052
+SMASH 69.3% 0.274 84.5% 0.150 67.7% 0.246
+DSTPP 0.5% 0.987 30.9% 0.691 32.3% 0.892
+ETAS consistently performs best across all datasets and tests.
+It achieves the highest pass rates and lowest KS statistics, indicating strong calibration and reliability.
+SMASH shows moderate performance, often outperforming DSTPP but trailing ETAS.
+Its results vary more across datasets and tests, with occasional strengths (e.g.
+in White for spatial KS).
+DSTPP generally performs worse, with lower pass rates and higher KS statistics, especially for the SCEDC and White datasets.
+However, it achieves relatively good spatial calibration in some cases (e.g., ComCat).
+Further analysis of the simulated forecasts in Appendix G provides insight into the consistency test results.
+Temporally, all models struggle to capture the highest-rate seismicity days, indicating substantial room for improvement in modelling the most hazardous periods.
+SMASH exhibits highly variable, spiky daily rate forecasts that result in frequent over- and underprediction, while DSTPP produces much smoother forecasts that systematically underestimate seismicity across both background and active periods.
+Spatially, ETAS assigns concentrated rates along known fault structures through its explicit clustering mechanism.
+In contrast, SMASH generates diffuse spatial forecasts with weak contrast between active and inactive regions, whereas DSTPP more accurately follows fault-aligned structure but often assigns uniformly low spatial rates, particularly in the SCEDC and White datasets.
+We were unable to apply the CSEP evaluation procedure for NSTPP, AutoSTPP and DeepSTPP, since the models are not explicitly formulated to be generative and therefore suffer from slow sampling (see details in Appendix D).
+This limitation significantly hinders their ability to be applied to real-time operational earthquake forecasting.
 5 Discussion and Conclusion
@@
 We establish two evaluation frameworks tailored to seismology: standard log-likelihood metrics and the generative consistency tests developed by the Collaboratory for the Study of Earthquake Predictability (CSEP), ensuring that successful models can be directly relevant to operational forecasting.
-In benchmarking five NPP models against ETAS, we found that none outperformed the baseline, suggesting current NPPs are not yet suitable for operational use.
-While NPPs often perform well during low-activity background periods, they struggle during active phases following large earthquakes.
-This may be due to their lack of explicit magnitude dependence — a key feature in ETAS — and their limited memory, as NPPs encode only a fixed-length history.
-In contrast, ETAS sums over all past events, allowing long-past earthquakes to influence future rates.
-Future NPPs should incorporate magnitude information and develop mechanisms for longer-term memory.
-EarthquakeNPP, available at https://github.com/ss15859/EarthquakeNPP/tree/main, provides a platform for future NPP developments to be benchmarked against these initial results.
+In benchmarking five neural point process (NPP) models against ETAS, we found that none outperformed the baseline, indicating that current NPP architectures are not yet suitable for operational earthquake forecasting.
+While several NPPs achieve competitive performance during low-activity background periods, they consistently struggle during highly active phases following large earthquakes.
+Our results highlight several concrete architectural and methodological gaps, which we summarise below as actionable directions for future NPP development.
+Action 1: Encode explicit magnitude dependence.
+ETAS explicitly encodes magnitude dependence, whereby larger earthquakes exponentially increase both the rate and spatial extent of subsequent seismicity.
+None of the benchmarked NPPs incorporate such explicit magnitude scaling, which limits their ability to capture the dominant influence of large events.
+Future NPP architectures could address this by introducing magnitude-aware design choices, such as hierarchical encodings that distinguish small and large events, magnitude-weighted attention mechanisms, or parameterisations aligned with the logarithmic frequency–magnitude scaling observed in seismicity (Richter, 1935).
+These approaches would allow NPPs to retain flexibility while incorporating structure that has proven critical for ETAS performance.
+Action 2: Design scalable long-term memory mechanisms.
+All evaluated NPPs truncate the conditioning history due to the computational cost of sequence encoders, with models such as DeepSTPP and AutoSTPP conditioning on as few as 20 past events.
+In contrast, ETAS integrates the full event history, allowing long-past earthquakes, including large or spatially distant events, to influence future rates.
+Designing NPPs with scalable long-term memory is therefore a critical avenue for improvement.
+Promising directions include sparse or dilated attention mechanisms (Child et al., 2019; Hassani and Shi, 2022) to reduce quadratic complexity, hierarchical or coarse-to-fine representations of earthquake histories (Yang et al., 2016), and explicit memory compression modules (Kim et al., 2023) that preserve the influence of distant but significant events.
+Advances in long-context modelling within the NLP literature suggest that such mechanisms are technically feasible (Liu et al., 2025) and may translate naturally to earthquake triggering dynamics.
+Action 3: Align generative training with operational evaluation.
+Third, our results reveal a mismatch between how generative NPPs are trained and how they are evaluated.
+Models such as SMASH and DSTPP are trained to predict or sample the next event, whereas CSEP consistency tests require simulating complete event sequences over fixed forecasting windows.
+This discrepancy helps explain why some generative models show reasonable short-term accuracy (Yuan et al., 2023; Li et al., 2023) but perform poorly in our multi-event simulation tests.
+Future generative NPPs may therefore benefit from training objectives that explicitly target long-horizon trajectory behaviour, for example by optimising multi-event simulation losses (e.g.
+Lüdke et al., 2024) or designing statistics (e.g.
+equation 13) aligned with the CSEP evaluations used in this study.
+Action 4: Incorporate empirically supported scaling laws.
+Finally, our results suggest that the complete removal of physically motivated structure may be counterproductive.
+While NPPs aim to move beyond parametric models, ETAS kernels encode power-law scaling relationships that are strongly supported by empirical seismology (Kagan, 1994; Felzer and Brodsky, 2006).
+Hybrid architectures that combine neural density estimation with ETAS-inspired power-law kernels or magnitude-dependent triggering functions may offer a productive middle ground, retaining empirical laws while allowing greater flexibility than purely parametric formulations.
+For example, replacing Gaussian spatial kernels in existing NPPs (e.g in DeepSTPP) with learned power-law forms could improve their ability to represent aftershock clustering without sacrificing expressiveness.
+EarthquakeNPP, available at https://github.com/ss15859/EarthquakeNPP, provides a platform for future NPP developments to be benchmarked against these initial results.
 The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions.
 Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
-Acknowledgments and Disclosure of Funding
+Acknowledgments
 This project is funded by Compass - Centre for Doctoral Training in Computational Statistics and Data Science (EPSRC Grant Ref EP/S023569/1).
 Compass is funded by United Kingdom Research and Innovation (UKRI) through the Engineering and Physical Sciences Research Council (EPSRC), https://www.ukri.org/councils/epsrc.
-This project also has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant 821115, Real-time earthquake rIsk reduction for a reSilient Europe (RISE), http://www.rise-eu.org) and by United States Geological Survey (USGS) EHP grant G24AP00059.
+This project also has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant 821115, Real-time earthquake rIsk reduction for a reSilient Europe (RISE), http://www.rise-eu.org) and by United States Geological Survey (USGS) EHP grants G24AP00059 and G25AP00379.
 Appendix A Earthquake Catalog Data
@@
 Below, we briefly overview the process of generating an earthquake catalog to illustrate the data quality issues.
-In the subsequent section, we review catalog incompleteness and its potential impact on the performance and evaluation of forecast models.
+In the subsequent section, we review catalog incompleteness and its potential impact on the performance and evaluation of forecasting models.
 Seismometers and Seismic Networks.
-A seismometer is an instrument that detects and records the vibrations caused by seismic waves [75, 71].
+A seismometer is an instrument that detects and records the vibrations caused by seismic waves (Stein and Wysession, 2009; Shearer, 2019).
 It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities.
-Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., [92] and references therein).
+Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., (Woessner et al., 2010) and references therein).
 High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
@@
 The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake.
-Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes [2].
-Recent algorithms, such as machine learning classifiers [e.g.
-98, 35] and template-matching [e.g.
-62], can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
+Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes (Allen, 1982).
+Recent algorithms, such as machine learning classifiers (e.g.
+Zhu and Beroza, 2019; Lapins et al., 2021) and template-matching (e.g.
+Ross et al., 2019), can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
 Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake.
 Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event.
-Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms [82, 37].
-Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist [e.g.
-92, and references therein].
+Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms (Thurber, 1985; Lomax et al., 2000).
+Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist (e.g.
+Woessner et al., 2010, and references therein).
 Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
-Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter [61].
+Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter (1935).
 The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers.
 This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons.
-This led to introduction of the moment magnitude scale (Mw) [22], which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
+This led to introduction of the moment magnitude scale (Mw) (Hanks and Kanamori, 1979), which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
 $M_{0}=\text{rigidity}\times\text{rupture area}\times\text{slip},$ (7)
@@
 Mw is determined seismologically via a spectral fitting process to the earthquake waveforms.
-In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales [e.g.
-26, and references therein].
+In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales (e.g.
+Herrmann and Marzocchi, 2021, and references therein).
 A.2 Earthquake Catalog Completeness
@@
 However, constructing a suitable retrospective forecasting experiment from this raw data requires appropriate pre-processing.
-This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ [e.g., 42, 43].
+This typically involves truncating the dataset above a magnitude threshold $M_{\text{cut}}$ and within a target spatial region to address incomplete data, known as catalog completeness $M_{c}$ (e.g., Mignan et al., 2011; Mignan and Woessner, 2012).
 There are several reasons why an earthquake may not be detected by a seismic network.
 Small events may be indistinguishable from noise at a single station, or insufficiently corroborated across multiple stations.
-Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high [34, 20].
+Another significant cause of missing events occurs during the aftershock sequence of large earthquakes, when the seismicty rate is high (Kagan and Knopoff, 1987; Hainzl, 2022).
 Human or algorithmic detection abilities are hampered when numerous events occur in quick succession, e.g.
@@
 Choosing a truncation threshold $M_{\text{cut}}$ that is too high removes usable data.
-Where NPPs have demonstrated an ability to perform well with incomplete data [77], typically a threshold below the completeness biases classical models such as ETAS [67].
-Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds [e.g.
-40, 77], which we also facilitate in our datasets.
-Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law [17], which states that the distribution of earthquake magnitudes follows an exponential probability density function
+Where NPPs have demonstrated an ability to perform well with incomplete data (Stockman et al., 2023), typically a threshold below the completeness biases classical models such as ETAS (Seif et al., 2017).
+Seismologists often investigate the biases of different magnitude thresholds by performing repeat forecasting experiments for different thresholds (e.g.
+Mancini et al., 2022; Stockman et al., 2023), which we also facilitate in our datasets.
+Typically $M_{c}$ is determined by comparing the raw earthquake catalog to the Gutenberg-Richter law (Gutenberg and Richter, 1936), which states that the distribution of earthquake magnitudes follows an exponential probability density function
 $f_{GR}(m)=\beta e^{-\beta(m-M_{c})}\ \ :m\geq M_{c}.$ (8)
 where $\beta$ is a rate parameter related to the b-value by $\beta=b\log 10$.
-Histogram-based approaches, such as the simple Maximum Curvature method [91] as well as many others [e.g.
-26, and references therein], identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
-In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ [e.g.
-65].
-During aftershock sequences, $M_{c}(t)$ can be very high [e.g., 1, 19] (See Figure 6a).
+Histogram-based approaches, such as the simple Maximum Curvature method (Wiemer and Wyss, 2000) as well as many others (e.g.
+Herrmann and Marzocchi, 2021, and references therein), identify the magnitude at which the observed catalog deviates from this law, indicating incompleteness (See Figure 6b).
+In practice, catalog completeness varies in both time and space $M_{c}(t,\mathbf{x})$ (e.g.
+Schorlemmer and Woessner, 2008).
+During aftershock sequences, $M_{c}(t)$ can be very high (e.g., Agnew, 2015; Hainzl, 2016b) (See Figure 6a).
 Thresholding at the maximum value might remove too much data.
-Instead, modelers either omit particularly incomplete periods during training and testing [32, 21], model the incompleteness itself [25, 89, 55, 18, 19, 45, 20], or accept known biases from disregarding this issue [73].
-Spatially, catalogs are less complete farther from the seismic network [42], so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
+Instead, modelers either omit particularly incomplete periods during training and testing (Kagan, 1991; Hainzl et al., 2008), model the incompleteness itself (Helmstetter et al., 2006; Werner et al., 2011; Omi et al., 2014; Hainzl, 2016a; b; Mizrahi et al., 2021; Hainzl, 2022), or accept known biases from disregarding this issue (Sornette and Werner, 2005).
+Spatially, catalogs are less complete farther from the seismic network (Mignan et al., 2011), so the spatial region can be constrained to remove outer, more incomplete areas (See Figure 6c).
 Figure 6: a) the June 10, 2016 Mw5.2 Borrego Springs earthquake and aftershocks, which occurred on the San Jacinto fault zone and is recorded in the White catalog.
@@
 Beyond the official EarthquakeNPP datasets, we include 3 further datasets that either provide additional scientific insight or continuity from previous benchmarking works.
-Synthetic ETAS Catalogs.
-We simulate a synthetic catalog using the ETAS model with parameters estimated from ComCat, at $M_{c}\ 2.5$, within the same California region.
-A second catalog emulates the time-varying data-missingness present in observational catalogs by removing events using the time-dependent formula from Page et al.
-[59],
-$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (9)
-where $M$ is the mainshock magnitude.
-Events below this threshold are removed using mainshocks of Mw 5.2 and above.
-The inclusion of these datasets allows us to test whether NPPs are inhibited by data missingness to the same extent that ETAS is.
-Deprecated Catalog of Japan.
-To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al.
-[5], however with a chronological train-test split and without removing any supposed outlier events.
-To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
 Table 3: Summary of additional datasets, including: magnitude threshold ($\mathbf{M}_{\mathbf{c}}$), number of training events, and number of testing events.
@@
 Japan_Deprecated 2.5 1990-1992-2007-2011-2020 22,213 15,368
+B.1 Synthetic ETAS Catalogs.
+We simulate a synthetic catalog using the ETAS model with parameters estimated from ComCat, at $M_{c}\ 2.5$, within the same California region.
+A second catalog emulates the time-varying data-missingness present in observational catalogs by removing events using the time-dependent formula from Page et al.
+(2016),
+$M_{c}(M,t)=M/2-0.25-\log_{10}(t),$ (9)
+where $M$ is the mainshock magnitude.
+Events below this threshold are removed using mainshocks of Mw 5.2 and above.
+The inclusion of these datasets allows us to test whether NPPs are inhibited by data missingness to the same extent that ETAS is.
+B.2 Deprecated Catalog of Japan.
+To provide continuity from the previous benchmarking for NPPs on earthquakes, we also provide results on the Japanese dataset from Chen et al.
+(2021), however with a chronological train-test split and without removing any supposed outlier events.
+To reflect our recommendation not to use this dataset in any future benchmarking following the dataset completeness issues mentioned above, we name this dataset Japan_Deprecated.
+We can use this corrected dataset to quantify the inflation of performance caused by the non-chronological training-validation-testing splitting in the Chen et al.
+(2021) dataset.
+Table 4 presents the information gain (difference in total log-likelihood, see section 2.1) relative to a Poisson process for the three NPP models across the two datasets.
+The dramatic drop in information gain highlights how the original data split and omission of the 2011 Tohoku earthquake inflates model performance.
+Table 4: NPP performance comparison on the Original Chen et al. (2021) dataset versus the Japan_Deprecated dataset. Values are reported in terms of information gain from a homogeneous Poisson process.
+Model Original Chen et al. (2021) dataset Japan_Deprecated
+NSTPP 11.26 1.99
+DeepSTPP 11.77 2.95
+AutoSTPP 12.16 2.76
+B.3 Likelihood Evaluation
 Figures 7 and 8 report the temporal and spatial log-likelihood scores of all the benchmarked models on additional datasets.
@@
 Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
-Appendix C Computational Efficiency
-C.1 Training
+Appendix C Effect of Training Window on ETAS Performance
+To verify that fitting ETAS on both the training and validation windows does not artificially improve its performance relative to the NPPs, we retrained ETAS using only the training window and re-evaluated its test log-likelihoods.
+As shown in Figures  9 & 10, the resulting log-likelihood scores are effectively unchanged across all EarthquakeNPP datasets.
+Figure 9: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
+Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
+ETAS (orange) is trained on both training and validation windows, whereas ETAS* (light blue) is trained only using the training window.
+Figure 10: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets.
+Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
+ETAS (orange) is trained on both training and validation windows, whereas ETAS* (light blue) is trained only using the training window.
+Appendix D Computational Efficiency
+D.1 Training
 Table 5 reports the training times for each model across all datasets.
 We ran all the NPP models using a HPC node with Nvidia Ampere GPU with 4x Nvidia A100 40GB SXM “Ampere” GPUs and AMD EPYC 7543P 32-Core Processor “Milan” CPU using torch==1.12.0 and cuda==11.3.
+[…]
 Dataset # Training Events ETAS Deep-STPP AutoSTPP NSTPP SMASH DSTPP Poisson
@@
 ETAS_incomplete 115,115 00:35:14 00:15:29 01:09:43 3 days, 11:39:51 1:44:34 1 day, 2:28:42 <1 second
-[…]
 ETAS training scales $\mathcal{O}(n^{2})$ with the total number of events, since for every event a contribution to the intensity function is computed from a summation over all previous events.
 This scaling, coupled with the lack of parallelization in the current implementation, results in long training times for larger datasets.
-Poorer scaling will likely hinder ETAS if dataset sizes continue to grow in the future [76].
+Poorer scaling will likely hinder ETAS if dataset sizes continue to grow in the future (Stockman et al., 2024).
 Encouragingly, both DeepSTPP and AutoSTPP are significantly faster to train due to GPU acceleration and their use of a sliding window of the most recent $k=20$ events.
 While exact complexity analyses are not provided in Zhou et al.
-[97] or Zhou and Yu [96], we can infer that DeepSTPP likely scales as $\mathcal{O}(kn)$ since it benefits from a closed-form expression for the likelihood.
+(2022) or Zhou and Yu (2024), we can infer that DeepSTPP likely scales as $\mathcal{O}(kn)$ since it benefits from a closed-form expression for the likelihood.
 AutoSTPP, though requiring automatic integration to compute the likelihood, still scales with $\mathcal{O}(kn)$ because the additional integration cost does not affect the overall scaling.
@@
 Unlike the sliding window mechanism used in DeepSTPP and AutoSTPP, NSTPP partitions the event sequence into fixed time intervals, leading to sequences that are much longer than the $k=20$ events used by the other models (as shown in Figure 11 of Chen et al.
-[5]).
+(2021)).
 Furthermore, solving an ODE for each event time adds a significant computational burden, even with the use of their faster attentive CNF architecture.
-Whilst SMASH and DSTPP are built on the same backbone architecture, SMASH is much quicker to train than DSTPP—even faster than ETAS.
+Whilst SMASH and DSTPP are built on the same backbone architecture, SMASH is much quicker to train than DSTPP, even faster than ETAS.
 This efficiency arises from its use of a single-step, normalization-free score-matching objective, which avoids the costly denoising and sampling loops required in diffusion-based training.
@@
 In contrast, DSTPP simulates a sequential generative process over hundreds of intermediate steps per sample, significantly increasing computation and memory costs.
-C.2 Simulation
-Real-time earthquake forecasting and CSEP model evaluation require simulating many repeat sequences (at least 10,000 for adequate distributional coverage) over the forecasting horizon.
-While ETAS training scales as $\mathcal{O}(n^{2})$ with the number of training events, its simulation scales more efficiently at $\mathcal{O}(n\log n)$.
-This improved scaling is due to its equivalent formulation as a Hawkes branching process (see Section 2.2).
-Both Deep-STPP and AutoSTPP are also based on Hawkes processes, which theoretically allows for fast simulation.
-However, as these models currently only have an intensity function implementation, simulating events would require a slower thinning procedure [52], limiting their simulation efficiency.
-NSTPP too has slow simulation, since it must solve an ODE to sample a new event.
-Due to the slow simulation of DeepSTPP, AutoSTPP and NSTPP, they are not evaluated using the CSEP generative evaluation metrics.
-Appendix D Analysis of Likelihood Scores
+D.2 Inference
+Whilst log-likelihood computation for ETAS, Poisson, DeepSTPP, AutoSTPP, and NSTPP is fast (<30<30 seconds per dataset), real-time earthquake forecasting and CSEP evaluation require simulating many repeated event sequences (at least 10,000) over a fixed forecasting horizon.
+While ETAS training scales as $\mathcal{O}(n^{2})$ with the number of training events, its simulation is considerably more efficient, scaling approximately as $\mathcal{O}(n\log n)$ due to its equivalent formulation as a Hawkes branching process (see Section 2.2), which enables immigration–birth sampling.
+Fast simulation is not currently feasible for several NPP architectures.
+NSTPP is not Hawkes-based and requires solving a neural ordinary differential equation to generate each new event, rendering simulation prohibitively slow even for small datasets.
+DeepSTPP and AutoSTPP, while inspired by Hawkes processes, employ non-stationary neural triggering kernels that depend on the full event history.
+As a result, standard immigration–birth sampling cannot be applied, since triggering relationships change after each event and new events cannot be generated independently or in parallel.
+Simulation via thinning is also problematic: AutoSTPP does not enforce monotonic or decaying kernels, meaning the conditional intensity $\lambda^{*}(t,\mathbf{x})$ cannot be safely upper-bounded, while DeepSTPP can in principle be simulated via thinning but is extremely slow in practice.
+For these reasons, DeepSTPP, AutoSTPP, and NSTPP are excluded from CSEP generative evaluation.
+For the models where large-scale simulation is tractable (ETAS, SMASH, and DSTPP), we report inference and simulation times in Table 6.
+Based on simulation times, DSTPP is fastest, followed by ETAS, with SMASH consistently the slowest across datasets, reflecting the fact that DSTPP samples events via a fixed-length diffusion process with closed-form updates, whereas SMASH relies on iterative Langevin dynamics requiring many gradient evaluations per event.
+Table 6: Simulation time for a batch of 100 repeated simulations across datasets.
+Reported times correspond to individual forecast days and are summarised as minimum, median, mean, and maximum over the testing window, reflecting variation in daily event counts rather than runtime stochasticity.
+Times are formatted as $H$:$M$:$S$.
+Dataset No. Days Daily Counts ETAS SMASH DSTPP
+(min / med / mean / max) (min / med / mean / max) (min / med / mean / max) (min / med / mean / max)
+ComCat 4764 0 / 3 / 4.59 / 1241 00:00:24 / 00:01:25 / 00:01:33 / 01:49:25 00:02:40 / 00:03:41 / 00:04:20 / 08:26:00 00:00:08 / 00:01:10 / 00:01:12 / 00:31:20
+SaltonSea 731 0 / 2 / 5.61 / 292 00:00:19 / 00:01:20 / 00:01:39 / 00:26:38 00:02:15 / 00:03:17 / 00:04:45 / 02:00:57 00:00:06 / 00:01:08 / 00:01:14 / 00:08:12
+SanJac 731 0 / 5 / 6.02 / 241 00:00:34 / 00:01:35 / 00:01:41 / 00:22:11 00:03:28 / 00:04:30 / 00:04:55 / 01:40:15 00:00:12 / 00:01:13 / 00:01:14 / 00:06:58
+SCEDC 2191 0 / 3 / 5.96 / 2134 00:00:23 / 00:01:25 / 00:01:40 / 03:07:19 00:02:40 / 00:03:41 / 00:04:54 / 14:28:21 00:00:07 / 00:01:10 / 00:01:14 / 00:53:05
+WHITE 1461 0 / 13 / 16.48 / 520 00:01:15 / 00:02:17 / 00:02:36 / 00:46:31 00:06:43 / 00:07:45 / 00:09:10 / 03:33:28 00:00:23 / 00:01:24 / 00:01:30 / 00:13:46
+Appendix E Analysis of Likelihood Scores
+E.1 Temporal Information Gain
+To better interpret the temporal likelihood results presented in Figure 2, we analyse how model performance evolves over time in each dataset in Figure 11, presenting the cumulative information gain (log-likelihood difference) of each NPP model over ETAS.
+In Figure 11(a) on the ComCat dataset, the largest decreases in relative performance occur during the two largest sequences in the testing window, namely the 2010 El Mayor Cucapah M7.2 and the 2019 Ridgecrest M7.1 earthquakes.
+ETAS performs strongly in these periods, likely because it incorporates magnitude scaling which enables it to model large aftershock cascades effectively.
+In contrast, the 2014 South Napa M6.0 and 2014 Offshore Eureka M6.8 do not produce such a marked drop in relative performance.
+This suggests that capturing the largest aftershock sequences is a key limitation of current NPP models, while during background periods NPPs perform relatively better due to their ability to capture non-stationarity, a property directly not modelled by ETAS.
+Despite the overall worse performance of NSTPP, the relative decrease during large events is not as sharp as the other models, suggesting superior performance.
+A similar pattern appears in Figure 11(b) for the SCEDC dataset.
+All models show a sharp reduction in performance during the 2019 Ridgecrest M7.1 sequence.
+During the quieter period leading up to Ridgecrest, DeepSTPP performs comparably to ETAS.
+This again suggests that the lack of explicit magnitude conditioning limits NPP performance during large sequences and highlights NPP models ability to capture background non-staionarities.
+Figures 11(c), 11(d) and 11(e) show results for the smaller regional datasets, SanJac, SaltonSea and White.
+These smaller magnitude, more regionally concentrated datasets, don’t display the same constrasting "background", "mainshock" behaviour present in ComCat and SCEDC.
+Although the overall performance of the NPP models is below that of ETAS, improvements occur during the 2017 Brawley swarm in the SaltonSea dataset, whereby DeepSTPP achieves higher temporal likelihood than ETAS.
+This behaviour aligns with the known difficulty ETAS has in modelling swarm-like sequences that are not initiated by a large mainshock.
 (a) ComCat
@@
 Figure 11: Cumulative information gain (IG) plots for the temporal performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
+E.2 Spatial Information Gain
+To better interpret the spatial likelihood results presented in Figure 3, we analyse how model performance evolves over space in each dataset in Figures 12, 13, visualising geographically where NPPs outperform or fall behind ETAS.
+Figures 12 and 13 show the spatial distribution of log-likelihood information gain across all datasets.
+Across both figures, a consistent pattern emerges.
+ETAS performs best in regions dominated by large, magnitude driven mainshock–aftershock sequences, while NPP performance degrades in the immediate vicinity of these events.
+In contrast, NPPs tend to perform more competitively in regions characterised by spatially complex or diffuse seismicity.
+In the larger regional catalogs shown in Figure 12, this distinction is most apparent in the ComCat dataset.
+Near the Ridgecrest and El Mayor–Cucapah sequences, NPPs exhibit reduced information gain relative to ETAS, consistent with the absence of explicit magnitude driven triggering.
+However, in the complex tectonic setting of the Mendocino Triple Junction, NPPs achieve comparatively strong spatial performance, with frequent positive information gain relative to ETAS.
+This region is characterised by interacting fault systems and persistent background activity rather than a single dominant mainshock, suggesting that NPPs are better suited to modelling such non-stationary and spatially heterogeneous seismic regimes.
+Figure 13 presents results for the smaller regional datasets San Jacinto, Salton Sea, and White.
+These regions are dominated by lower magnitude seismicity and swarm like behaviour, and here NPPs again perform more competitively relative to ETAS.
+AutoSTPP assigns probability more smoothly along fault structures similarly to ETAS, resulting in information gain values concentrated near zero, while DeepSTPP and NSTPP show more heterogeneous behaviour.
+NSTPP produces more extreme and spikier information gain values that are spatially scattered, indicating less precise spatial localisation despite occasional high likelihood assignments.
+Instability during training led to the drastic underperformance of NSTPP on the White dataset and consequent low likelihood scores distributed across the entire region.
+Overall, the spatial results mirror the temporal analysis.
+Current NPP architectures struggle most in mainshock dominated regimes but show clear promise in modelling spatially complex background seismicity and swarm driven activity, motivating future work on incorporating large magnitude triggering while preserving this flexibility.
 (a) ComCat
 (b) SCEDC
-(c) SanJac
-(d) SaltonSea
-(e) WHITE
-Figure 10: Cumulative information gain (IG) plots for the spatial performance of all the NPP models with respect to ETAS on a) ComCat,b) SCEDC, c) QTM_San_Jac, d) QTM_Salton_Sea, e) White.
-Appendix E CSEP Consistency Tests
-E.1 Number (Temporal) Test
+Figure 12: Spatial information gain per event, for NPP models relative to ETAS on (a) ComCat and (b) SCEDC.
+Scatter points correspond to the geographical location of the forecasted event, coloured by the value of the information gain over the ETAS model (green positive, red negative).
+For each model an inset plot displays the distribution of the spatial information gains for all events in the testing period.
+(a) SanJac
+(b) SaltonSea
+(c) White
+Figure 13: Spatial information gain per event, for NPP models relative to ETAS on (a) SanJac, (b) SaltonSea, and (c) White.
+Scatter points correspond to the geographical location of the forecasted event, coloured by the value of the information gain over the ETAS model (green positive, red negative).
+For each model an inset plot displays the distribution of the spatial information gains for all events in the testing period.
+Appendix F CSEP Consistency Tests
+F.1 Number (Temporal) Test
 The number test evaluates the temporal component of the forecast by checking the consistency of the forecasted number of events, $N$ with those observed in the forecast horizon, $N_{\text{obs}}$.
@@
 $\delta_{2}$ $=\mathbb{P}(N\leq N_{\text{obs}})=F_{N}(N_{\text{obs}}).$ (11)
-E.2 Pseudo-Likelihood Test
+F.2 Pseudo-Likelihood Test
 The pseudo-likelihood test evaluates the compatibility of a forecast with an observed catalog using an approximation to the space-time point process likelihood.
@@
 $\gamma_{L}=\mathbb{P}(\hat{L}_{j}\leq\hat{L}_{\text{obs}})=F_{L}(\hat{L}_{\text{obs}}).$ (14)
-E.3 Spatial Test
+F.3 Spatial Test
 To evaluate the spatial component of the forecast, a test statistic aggregates the forecasted rates of earthquakes over a regular grid,
@@
 The grid is constructed from $\{0.1\degree,0.05\degree,0.01\degree\}$ squares for ComCat, SCEDC and {QTM_Salton_Sea, QTM_SanJac, White} respectively.
-E.4 Magnitude Test
+F.4 Magnitude Test
 To evaluate the earthquake magnitude component of the forecast, a test statistic compares the histogram of a forecast’s magnitudes $\Lambda^{(m)}$, against the mean histogram over all forecasts $\bar{\Lambda}^{(m)}$,
@@
 Histogram bins of size $\delta_{m}=0.1$ are used across all datasets.
+Although SMASH is, in principle, capable of modelling earthquake magnitudes, we restrict it to rate forecasting in this benchmark, as extending it to fine-grained magnitude prediction led to a deterioration in spatio-temporal performance metrics during training.
 (a) Number test
@@
 The test fails if the observed statistic falls within the rejection region (red), defined by the $0.05$ and $0.95$ quantiles of the distribution.
-E.5 Evaluating Multiple Forecasting Periods
+F.5 Evaluating Multiple Forecasting Periods
 Savran et al.
-[63] describe how to assess a model’s performance across the multiple days in the testing period (Figure 12).
-By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator [15].
+(2020) describe how to assess a model’s performance across the multiple days in the testing period (Figure 15).
+By construction, quantile scores over multiple periods should be uniformly distributed if the model is the data generator (Gneiting and Katzfuss, 2014).
 Therefore comparing quantile scores against standard uniform quantiles (y = x), highlights discrepancies between the observed data and the forecast.
@@
 Red bars indicate failed forecasts, and the Kolmogorov–Smirnov (KS) statistic quantifies deviation from uniformity.
-Appendix F Further Dataset Figures
-F.1 ComCat
+Appendix G Analysis of CSEP Tests
+G.1 Temporal
+To further interpret the CSEP consistency test results reported in Table 2, Figures 1618 show the daily event count forecasts produced by ETAS, SMASH, and DSTPP across all EarthquakeNPP datasets.
+These plots reveal systematic differences in how the models capture both background seismicity and sudden increases in earthquake rate.
+Across all datasets, ETAS provides the most consistent forecasts of daily event counts.
+It captures low-activity days well and responds more effectively than the NPP-based models to increases in seismicity rate, although it still underpredicts the largest rate excursions associated with major earthquake sequences.
+This behaviour is particularly evident in the ComCat and SCEDC datasets during the 2010 El Mayor–Cucapah and 2019 Ridgecrest sequences (Figures 16(a) and 16(b)), and explains ETAS’s consistently high consistency test pass rates.
+SMASH exhibits highly variable daily rate estimates across all regions.
+While this variability occasionally allows it to match periods of elevated seismicity, such as during El Mayor–Cucapah in ComCat, it more often leads to pronounced over and under prediction.
+This spiky behaviour results in a substantial number of failed consistency tests.
+DSTPP produces much smoother daily rate forecasts than SMASH, but this comes at the cost of systematic underprediction.
+Across all datasets, DSTPP underestimates both background seismicity and elevated rate periods, with the effect becoming especially severe in SCEDC and White (Figures 16(b) and 18).
+This persistent bias explains its low consistency test pass rates.
+Performance differences across datasets largely reflect the dominant seismic regime.
+In smaller regions such as San Jacinto and Salton Sea (Figures 17(a) and 17(b)), all models perform more competitively due to the absence of large mainshock-driven rate increases.
+However, even in these settings, SMASH remains overly variable and DSTPP continues to underestimate daily rates.
+(a) ComCat
+(b) SCEDC
+Figure 16: Daily number forecasts from ETAS, SMASH, and DSTPP over the full testing period for (a) ComCat and (b) SCEDC.
+Vertical lines show the forecasted daily distributions of earthquake counts.
+Green lines indicate days where the observed count falls within the 95% forecast interval, while red lines indicate failures.
+Observed daily counts are shown as dots, with marker size proportional to the number of earthquakes and colour indicating pass (green) or fail (red).
+(a) SanJac
+(b) SaltonSea
+Figure 17: Daily number forecasts from ETAS, SMASH, and DSTPP over the full testing period for (a) SanJac and (b) SaltonSea.
+Vertical lines show the forecasted daily distributions of earthquake counts.
+Green lines indicate days where the observed count falls within the 95% forecast interval, while red lines indicate failures.
+Observed daily counts are shown as dots, with marker size proportional to the number of earthquakes and colour indicating pass (green) or fail (red).
+Figure 18: Daily number forecasts from ETAS, SMASH, and DSTPP over the full testing period for White.
+Vertical lines show the forecasted daily distributions of earthquake counts.
+Green lines indicate days where the observed count falls within the 95% forecast interval, while red lines indicate failures.
+Observed daily counts are shown as dots, with marker size proportional to the number of earthquakes and colour indicating pass (green) or fail (red).
+G.2 Spatial
+Figures 19 and 20 show aggregated spatial forecasts over the entire testing period for ETAS, SMASH, and DSTPP across all EarthquakeNPP datasets.
+These plots summarise how each model distributes seismicity rate spatially when forecasts are integrated over time.
+While this provides a useful overview of long-term spatial structure, it does not distinguish whether high rates are forecast before events occur.
+Across all datasets, ETAS produces the most spatially concentrated forecasts, with high rates aligned along known fault structures and low rates assigned to seismically inactive regions.
+This apparent spatial precision arises largely from ETAS’s explicit modelling of clustering, whereby elevated rates are assigned in the vicinity of earthquakes that have already occurred.
+This behaviour is particularly evident in the ComCat and SCEDC datasets (Figures 19(a) and 19(b)), where ETAS reproduces the large-scale fault network and major clusters associated with past seismicity.
+In contrast, SMASH consistently produces spatially diffuse forecasts.
+While it sometimes captures regions of elevated activity, its rate is spread broadly across each domain, leading to weaker spatial contrast between active and inactive regions.
+This effect is most pronounced in ComCat, where SMASH concentrates strongly around the southern end of the domain near the 2010 El Mayor–Cucapah sequence, while remaining diffuse elsewhere.
+Similar behaviour is observed in SCEDC, where SMASH is again dominated by the El Mayor region and assigns comparatively high rates across large areas of the domain.
+DSTPP generally produces smoother spatial forecasts than ETAS but with greater structure than SMASH.
+In ComCat, San Jacinto, and Salton Sea (Figures 19(a), 20(a), and 20(b)), DSTPP follows the main fault-aligned clustering while attenuating sharp spatial contrasts.
+However, in SCEDC and White, DSTPP substantially underestimates the overall seismicity rate, resulting in uniformly low spatial forecasts and visible boundary effects, particularly near the edges of the domains (Figures 19(b) and 20(c)).
+(a) ComCat
+(b) SCEDC
+Figure 19: Spatial forecasts from ETAS, SMASH and DSTPP, aggregated across the entire testing period of (a) ComCat and (b) SCEDC.
+For each day in the testing period, every model simulates 10,000 repeated earthquake catalogs within the boundary region.
+All simulated catalogs are aggregated across the entire testing period and the expected earthquake rates are plotted within each grid cell.
+Observed earthquakes are overlaid with marker size proportional to magnitude.
+(a) SanJac
+(b) SaltonSea
+(c) White
+Figure 20: Spatial forecasts from ETAS, SMASH and DSTPP, aggregated across the entire testing period of (a) SanJac, (b) SaltonSea, and (c) White.
+For each day in the testing period, every model simulates 10,000 repeated earthquake catalogs within the boundary region.
+All simulated catalogs are aggregated across the entire testing period and the expected earthquake rates are plotted within each grid cell.
+Observed earthquakes are overlaid with marker size proportional to magnitude.
+Appendix H Further Dataset Figures
+H.1 ComCat
 Figure 21: Times and magnitudes of events in the ComCat dataset (with key events labeled).
@@
 Figure 22: Locations of events in the ComCat dataset, labeled by their partition into auxiliary, training, validation and testing periods.
-F.2 SCEDC
+H.2 SCEDC
 Figure 23: Times and magnitudes of events in the SCEDC dataset (with key events labeled).
@@
 Figure 24: Locations of events in the SCEDC dataset, labeled by their partition into auxiliary, training, validation and testing periods.
-F.3 White
+H.3 White
 Figure 25: Times and magnitudes of events in the White dataset (with key events labeled).
@@
 Figure 26: Locations of events in the White dataset, labeled by their partition into auxiliary, training, validation and testing periods.
-F.4 QTM_SanJac
+H.4 QTM_SanJac
 Figure 27: Times and magnitudes of events in the QTM_SanJac dataset.
@@
 Auxiliary, training, validation and testing periods are indicated by colour and a further cumulative count of events is indicated in red.
-F.5 QTM_SaltonSea
+H.5 QTM_SaltonSea
 Figure 28: Times and magnitudes of events in the QTM_SaltonSea dataset.
@@
 Figure 29: Locations of events in the QTM_SanJac and QTM_SaltonSea datasets, labeled by their partition into auxiliary, training, validation and testing periods.
-Appendix G Error Distributions & Next-event metrics
+Appendix I Error Distributions & Next-event metrics
 Figure 30: The distribution of errors ($Y_{\text{obs}}-Y_{\text{pred}}$) for the Normal(0,1)(0,1), Exponential(1)(1), and Pareto(2)(2) distributions.